### v2

# Navier-Stokes Equations with Fractional Dissipation and Associated Doubly Stochastic Yule Cascades

Abstract
We identify a probabilistic structure known as self-similar doubly stochastic Yule (DSY) cascade associated with the deterministic Navier-Stokes equations (NSE) in $\mathbb{R}^{d}$ with fractional dissipation $(-\Delta)^{\gamma}$. Interestingly, such a structure is well-defined only in the scaling-supercritical regime $\gamma\in(\frac{1}{2},\frac{d+2}{4})$. We then characterize parametric regions of $(d,\gamma)$ that correspond to the non-explosive, explosive, and hyperexplosive behavior of the DSY cascade. Stochastic solution processes are constructed recursively, and their expectations yield solutions to the fractional NSE whenever these expectations exist. Explosion and geometric properties of the DSY cascade are then exploited to establish non-uniqueness and finite-time blowup results for a scalar partial differential equation associated with the fractional NSE thanks to a majorization principle for stochastic solution processes. In the special case $d=2$, we derive a closed form for the solution process and prove the finite-time loss of integrability of the solution process for sufficiently large initial data. Notably, this lack of integrability does not necessarily imply finite-time blowup of solutions to the fractional NSE. Indeed, for vortex-flow initial data, we show that the solution can be continued beyond the time of integrability breakdown by employing a notion of averaging.
1 Introduction
One way to approach the unresolved well-posedness issues of the incompressible 3D Navier–Stokes equations is through their natural scaling properties; see (2.1). This perspective highlights, for example, the existence of a scaling gap between quantities that can be controlled a-priori and the functional bounds that are critical or supercritical in nature; [8, 29, 38, 10].
Another manifestation of this scaling issue arises when considering the Navier–Stokes equations on $\mathbb{R}^{d}$, $d\geq 2$, with fractional dissipation $\gamma>0$, given by the Laplacian raised to a fractional power; see $\textup{(NSE)}_{d,\gamma}$. On and above the critical line
$\gamma=\frac{d+2}{4},$
there are a number of global-in-time well-posedness and regularity results, including the classical 2D Navier–Stokes equation; [27, 12, 26, 38, 33, 3, 23, 11].
In contrast, below the critical line, that is, in the supercritical regime $\gamma<(d+2)/4$, the questions of global well-posedness and regularity remain largely open, with some evidence suggesting finite-time blow-up and non-uniqueness of solutions [13, 1, 21]. This supercritical regime includes the physically relevant 3D case with $\gamma=1$. Incidentally, the critical case $\gamma=(d+2)/4$ corresponds precisely to the situation in which the energy, given by the $L^{2}$-norm of the solution, is scaling-invariant; see Section 2, and in particular (2.2).
While the most profound example relating partial differential equations and stochastic processes in mathematical physics is no doubt that of the heat/diffusion equation and Brownian motion, other somewhat surprising illustrations are known in other contexts, including KPP and wave equations [31, 9, 4], discounted Riccati-type equations [2, 20, 17, 19], and incompressible Navier-Stokes equations [28]. In this paper we investigate how the interplay between dimension, dissipation, and criticality in the fractional Navier–Stokes problem on $\mathbb{R}^{d}$ can be understood through a naturally associated stochastic structure originating with LeJan and Sznitman [28], where a scaling kernel $h(\xi)$ is naturally associated with the Fourier transformed equation and wave number $\xi$. As shown in Section 2, this structure arises directly from the natural rescaling of the Fourier-transformed fractional Navier–Stokes equations in the mild form; see (2.6)–(2.12).
A key element is the scaling kernel
$h(\xi)=c_{d,\gamma}|\xi|^{2\gamma-d-1},$
which can be derived by rescaling $\widehat{u_{\lambda}}$ with $\lambda=1/|\xi|$. Thus, the exponent $2\gamma-d-1$ is directly linked to the natural scaling (2.1); see [14] for detailed calculations in the case $d=3$, $\gamma=1$. Remarkably, the corresponding integral kernel $H$ given by (2.12) is a probability kernel yielding a stochastic representation of solutions precisely in the supercritical regime $1/2<\gamma<(d+2)/4$; here ”criticality” refers to the scaling invariance of the associated energy functional (2.2). Indeed, in this regime, the resulting equation (nFNS) can be interpreted as a mean flow equation associated with a stochastic structure that can be described in terms of two random fields:
(i) A doubly-stochastic Yule (DSY) random field $\mathbb{Y}$, indexed by the vertices of a labeled binary tree. The tree-indexed labels form a Markov process when restricted to any fixed ray emanating from the root, as discussed in Section 3, and
(ii) a stochastic solution process $\bf{X}$}, defined via a binary recursion on the evolution of $\mathbb{Y}$ on a common probability space $(\Omega,\mathscr{F},P)$; see Section 2.
The expected value of $\bf{X}$}, when it exists, satisfies the associated mean field equation, which provides the framework for studying various regularity questions for the mean flow, such as uniqueness versus non-uniqueness, finite-time blow-up, and related phenomena.
This framework leads to a natural unified approach to the incompressible fractional Navier–Stokes equations in $\mathbb{R}^{d}$, within the scaling-invariant functional setting $u\in L^{\infty}([0,T),PM^{a}),$ where $PM^{a}$ denotes the space of pseudo-measures defined in (2.3), with $a=d+1-2\gamma$.
Well-posedness of the three-dimensional Navier–Stokes equations in such settings was first studied in [28] using a closely related probabilistic approach, and later in [29] via analytical methods, establishing global existence of mild solutions with small initial data. More recently, [23] proved a number of local and global regularity results for mild solutions with initial data $u_{0}\in PM^{a,b}:=PM^{a}+PM^{b}$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$.
In this paper, we look at the connection between the dimension and dissipation and the nature of criticality of the fractional Navier-Stokes problem in $\mathbb{R}^{d}$ through the prism of this naturally associated stochastic structure. In the original approach, [28] introduced a coin-tossing device to eliminate the possibility of there being infinitely many branchings of the cascade in finite time, to be referred to as (stochastic) explosion. This coin tossing device was removed in [14], opening the possibility of using stochastic explosion as a mechanism for establishing non-uniqueness of solutions; [20, 19, 18].
Our goal is to focus on phase transitions of the stochastic structures $\bf{X}$} and $\mathbb{Y}$ depending on parameters $d$ and $\gamma$, and to explore their possible connections to regularity problems; see, for example, [7, 22, 14]. Several key elements of the analysis include: (a) The stochastic geometry underlying the representation of the solution process $\bf{X}$}; (b) Stochastic symmetries of both $\mathbb{Y}$ and $\bf{X}$} under groups of tree automorphisms; (c) Critical parameters governing stochastic explosion and hyperexplosion events for $\mathbb{Y}$, see Definitions 2 and 2; and (d) Integrability of the (minimal) solution process, while exploiting symmetries to continue solutions beyond the existence of the mean flow.
The main takeaways of this research are twofold.
First, we identify parametric regions in terms of $d$ and $\gamma$ where the DSY cascade $\mathbb{Y}$ exhibits either explosive or non-explosive behavior; see Section 3.2 and Figure 6. The numerical evidence shows existence of hyperexplosive regions for any dimension $d$, based on the new criteria, see Section 3.2, that hold for general self-similar DSY cascades as defined in [16]. This opens the possibility in the future to expand the probabilistic approach developed for the $\alpha$-Riccati equation, [20, 19], to the fractional Navier-Stokes equations. The former relies heavily on stochastic hyperexplosion, but for the fractional Navier–Stokes equation itself, stochastic explosion/hyperexplosion alone does not qualitatively strengthen existing existence and uniqueness results for mild solutions; see Section 4 and the subsequent remark. However, in the scalar case of the Montgomery–Smith equation, introduced in [32], and its fractional version (MS)-(FMS), the geometric structure of $\bf{X}$} prevents cancellations, so that explosive regimes naturally lead to non-uniqueness of solutions; see Section 4. Interestingly, in both the explosive and non-explosive cases, the mean flows $\mathbb{E}({\bf{X}})$ exhibit finite-time blow-up; [18]. Although the arguments in [18] concern a different kernel, in the specific case $d=3$, $\gamma=1$, they can be readily adapted to the fractional Montgomery–Smith equation on $\mathbb{R}^{d}$ across the entire supercritical range $\frac{1}{2}<\gamma<\frac{d+2}{4}.$
Perhaps counterintuitively, the explosive and especially hyperexplosive regimes for $\mathbb{Y}$ occur near the critical line $\gamma=(d+2)/4$ (see Figure 6), where existence and uniqueness results are more plausible and partial regularity theory is available (see, e.g., [11]). By contrast, the non-explosive regimes lie deeper in the supercritical region, where well-posedness results are expected to be weaker. This highlights the limitations of interpreting stochastic explosion of the DSY cascade as a direct metaphor for ill-posedness.
Another main takeaway of this research is that in the two-dimensional case the geometric structure of the bilinear term in the Navier–Stokes nonlinearity can be exploited to obtain a closed-form expression for the minimal solution process $\bf{X}$} (see Section 5). This permits one to establish symmetry conditions on the initial data which, together with the symmetries of the DSY cascade $\mathbb{Y}$, make it possible to use a modified notion of averaging to extend solutions beyond the critical time $t_{c}$ that marks the point at which $\bf{X}$} ceases to be integrable; see Section 5. We further show that in the case $d=2$, the critical time $t_{c}$ is finite when the initial data are sufficiently large; see (5).
While it is classical that rotationally invariant initial data in 2D lead to global solutions via cancellations in the nonlinearity, the result presented here is the first to reveal this cancellation directly at the level of the stochastic cascades. This perspective allows us to define the “mean evolution” of $\bf{X}$} beyond the integrability threshold $t_{c}$, thereby opening a pathway to bypass the limitations imposed by the use of majorizing principles (see Section 4), which have played a central role in earlier existence proofs [28, 18].
Nevertheless, a key challenge remains: There is a discrepancy between the loss of integrability of $\bf{X}$} (expressed by the finiteness of $t_{c}$) and the blow-up of the mean-flow equation. To address this gap requires development of a more general framework for continuing the evolution past $t_{c}$, perhaps by introducing new notions of averaging for $\bf{X}$} that exploit cancellations arising from symmetries in both the solution process $\bf{X}$} and the DSY cascade $\mathbb{Y}$. An illustrative result along these lines is provided in Section 5.
The paper is organized as follows. In Section 2 the incompressible fractional Navier-Stokes equations are introduced, both as a PDE and in terms of the stochastic structures: the DSY $\mathbb{Y}$ and the (minimal) solution process $\bf{X}$}. Their relationship via the mean flow equation is established in Section 2, where the notion of a minimal solution process and integrable solution process naturally appear. The latter notion of minimality involves a notion of stochastic explosion that is defined here as well. The general concept of solution process stands apart from that of an integrable solution process and is defined via a binary stochastic recursion that transforms initial data in accordance with the ‘bilinear product’ intrinsic to the incompressible mean flow. The stochastic/geometric structure of the DSY $\mathbb{Y}$ is analyzed in Section 3. In particular, stochastic explosion/non-explosion events are characterized in terms of the dimension $d$ and the fractional exponent parameters $\gamma$. In Section 4 the Montgomery-Smith model is introduced as a tool for obtaining estimates on the critical time governing integrability of the solution process via a comparison method, more generally developed in [18] for the incompressible Navier-Stokes equation. Section 5 involves a further specialization to $d=2$ dimensions, where the stochastic geometry and symmetries underlying the solution process $\bf{X}$} are significantly simplified, while preserving the key elements and critical phenomena noted above. The role of stochastic ‘cancellations’ due to symmetries in the initial data and stochastic solution process are a special focus of this paper, and are shown to lead simply to a closed form solution defined over the critical range of integrability. Interestingly, it is also shown that this locally defined solution extends to a global solution.
2 Navier-Stokes equations with Fractional Dissipation
Consider the initial value problem of the $(d,\gamma)$-fractional Navier-Stokes equations on $\mathbb{R}^{d}$ ($d\geq 2$)
$u_{t}+(-\Delta)^{\gamma}u+u\nabla u+\nabla p=0,\ \text{div}\,u=0,\ u(\cdot,0)=u_{0},$ $\textup{(NSE)}_{d,\gamma}$
where $u=u(x,t)$, $x\in\mathbb{R}^{d}$, $t\geq 0$. A natural scaling that preserves the system is
$u_{\lambda}(x,t)=\lambda^{2\gamma-1}u(\lambda x,\lambda^{2\gamma}t),\ \ p_{\lambda}(x,t)=\lambda^{4\gamma-2}p(\lambda x,\lambda^{2\gamma}t),\ \ u_{0\lambda}(x)=\lambda^{2\gamma-1}u_{0}(\lambda x).\ \$(2.1)
An apriori bound of $\textup{(NSE)}_{d,\gamma}$ is $E[u](t)\leq\frac{1}{2}\int_{\mathbb{R}^{d}}|u_{0}|^{2}dx$ for all $t\geq 0$, where
$E[u](t)=\frac{1}{2}\int_{\mathbb{R}^{d}}|u(x,t)|^{2}dx+\int_{0}^{t}\int_{\mathbb{R}^{d}}|(-\Delta)^{\gamma/2}u(x,s)|^{2}dxds$ (2.2)
is the energy. Under the natural scaling (2.1), the energy is scaled as $E[u_{\lambda}](t)=\lambda^{4\gamma-(d+2)}E[u](\lambda^{2\gamma}t)$. The system (NSE)d,γ is called critical if $\gamma=\frac{d+2}{4}$ (i.e. if the energy is scale-invariant), supercritical if $\gamma<\frac{d+2}{4}$, and subcritical if $\gamma>\frac{d+2}{4}$. The system is called hypodissipative for $\gamma<1$ and hyperdissipative for $\gamma>1$. Global regularity is known for the critical and subcritical regime $\gamma\geq\frac{d+2}{4}$ (see e.g. [26, 38, 33]). Local and global regularity holds for mild solutions for any $\gamma\in(\frac{1}{2},\frac{d+2}{4})$ and $u_{0}\in PM^{a,b}:=PM^{a}+PM^{b}$ where $PM^{a}$ is the space of pseudomeasure-type [23]
$PM^{a}=\{v\in\mathscr{S}^{\prime}(\mathbb{R}^{d}):\hat{v}\in L^{1}_{\rm loc}(\mathbb{R}^{d}),\,\text{esssup}\,|\xi|^{a}|\hat{v}(\xi)|<\infty\}.$ (2.3)
Local and global regularity of mild solutions in the hyperdissipative regime $\gamma\geq 1$ for initial data $u_{0}$ in supercritical Besov spaces and Triebel-Lizorkin spaces are obtained in [3].
The case of particular interest is $d=3$, in which partial regularity is known for the supercritical hyperdissipative regime $1<\gamma<5/4$ [26]. If $u_{0}\in H^{\delta}$ for some arbitrary $\delta>0$, then global regularity holds for any $\gamma\in[\frac{5}{4}-\epsilon,\frac{5}{4}]$ where $\epsilon>0$ depends on $\|u_{0}\|_{H^{\delta}}$ [11]. Finite-time blowup of solutions to a forced system is obtained in [13] for the hypodissipative regime $0\leq\gamma<\frac{2}{9}(22-8\sqrt{7})$.
$\textup{(NSE)}_{d,\gamma}$ has an integral form
$u(t)=e^{-t(-\Delta)^{\gamma}}u_{0}-\int_{0}^{t}e^{(s-t)(-\Delta)^{\gamma}}\mathbf{P}\text{div}(u\otimes u)ds$ (2.4)
where $\mathbf{P}$ denotes the Leray projection onto the divergence-free vector fields. Let $\hat{u}$ be the Fourier transform of $u$ in the spatial variable
$\hat{u}(\xi,t)=\mathscr{F}\{u(\cdot,t)\}(\xi)=c_{0}\int_{\mathbb{R}^{d}}u(x,t)e^{-ix\cdot\xi}dx,\ \ \ \text{where}\ c_{0}=(2\pi)^{-d/2}.$ (2.5)
Taking the Fourier transform of (2.4) and using the convolution identity $\mathscr{F}\{fg\}=c_{0}\mathscr{F}\{f\}*\mathscr{F}\{g\}$ (see [24]*p. 284), one has
$\hat{u}(\xi,t)=\hat{u}_{0}(\xi)e^{-|\xi|^{2\gamma}t}+c_{0}\int_{0}^{t}|\xi|e^{-|\xi|^{2\gamma}s}\int_{\mathbb{R}^{d}}\hat{u}(\eta,t-s)\otimes_{\xi}\hat{u}(\xi-\eta,t-s)d\eta ds.$ (2.6)
Here,
$a\otimes_{\xi}b=-i(e_{\xi}\cdot b)\pi_{\xi^{\perp}}a$ (2.7)
where $e_{\xi}=\xi/|\xi|$ is the unit vector in $\xi$-direction and $\pi_{\xi^{\perp}}$ is the perpendicular projection onto the hyperplane orthogonal to $\xi$. As remarked in [18], the vector product $\otimes$ in (nFNS) can be replaced by a symmetrized vector product $\odot$ defined by
$a\odot_{\xi}b=\frac{1}{2}(a\otimes_{\xi}b+b\otimes_{\xi}a)=-\frac{i}{2}((e_{\xi}\cdot b)\pi_{\xi^{\perp}}a+(e_{\xi}\cdot a)\pi_{\xi^{\perp}}b).$ (2.8)
One can rewrite (2.6) as
$\hat{u}(\xi,t)=\hat{u}_{0}(\xi)e^{-|\xi|^{2\gamma}t}+c_{0}\int_{0}^{t}|\xi|e^{-|\xi|^{2\gamma}s}\int_{\mathbb{R}^{d}}\hat{u}(\eta,t-s)\odot_{\xi}\hat{u}(\xi-\eta,t-s)d\eta ds.$ (2.9)
Let
$h(\xi)=c_{d,\gamma}|\xi|^{2\gamma-d-1},~~~{\textrm{where}}~~~{{c}_{d,\gamma}}={{\pi}^{-d/2}}\frac{\Gamma(2\gamma-1)\Gamma{{\left(\frac{d+1-2\gamma}{2}\right)}^{2}}}{\Gamma\left(\frac{d+2-4\gamma}{2}\right)\Gamma{{\left(\frac{2\gamma-1}{2}\right)}^{2}}}.$ (2.10)
For $\gamma\in(\frac{1}{2},\frac{d+2}{4})$, one can infer from the Fourier transform of radial functions [36, p. 205]
$\mathscr{F}(|x|^{-\alpha})=2^{\frac{d}{2}-\alpha}\frac{\Gamma(\frac{d-\alpha}{2})}{\Gamma(\frac{\alpha}{2})}|\xi|^{\alpha-d},\ \ \ \forall\,\alpha\in(0,d)$
that
$|\xi|^{2\gamma-1}h(\xi)=h*h(\xi).$ (2.11)
The function $h$ defined by (2.10) is known as a standard majorizing kernel [6]. The normalized Fourier transform $\chi(\xi,t)=c_{0}\frac{\hat{u}(\xi,t)}{h(\xi)}$ satisfies
$\chi(\xi,t)=\chi_{0}(\xi)e^{-|\xi|^{2\gamma}t}+\int_{0}^{t}|\xi|^{2\gamma}e^{-|\xi|^{2\gamma}s}\int_{\mathbb{R}^{d}}\chi(\eta,t-s)\otimes_{\xi}\chi(\xi-\eta,t-s)H(\eta|\xi)d\eta ds$ (nFNS)
where $\chi_{0}(\xi)=\chi(\xi,0)$,
$H(\eta|\xi)=\frac{h(\eta)h(\xi-\eta)}{|\xi|^{2\gamma-1}h(\xi)},$ (2.12)
The divergence-free constraint on the initial condition $u_{0}$ becomes an orthogonality constraint on $\chi_{0}$:
[…]
At this point, the solution $\chi$ to (nFNS) can be interpreted probabilistically as the expected value of a “solution process” $\mathbf{X}$ defined implicitly by the recursive relation
[…]
Here, $Y$ is an exponentially distributed random variable with mean $|\xi|^{-2\gamma}$; $W_{1}\in\mathbb{R}^{d}$ is a random variable independent of $Y$ with probability density $H(\cdot|\xi)$; $W_{2}=\xi-W_{1}$; and $\mathbf{X}^{(1)}$ and $\mathbf{X}^{(2)}$ are, conditionally given $Y$, two independent copies of $\mathbf{X}$. We state the result precisely as follows.
As usual, we use the Lebesgue measure on $\mathbb{R}^{k}$ and $\mathbb{C}^{k}$, $k\in\mathbb{N}$. A subset of a measurable set is said to have full measure if the complement has measure zero.
Definition 2.1 (Solution Process).
For a given measurable function $\chi_{0}:\mathbb{R}^{d}\to\mathbb{C}^{d}$, a stochastic process process $\mathbf{X}$ satisfying (2.14) in distribution, is referred to as a solution process. If $\mathbb{E}_{\xi}|{\mathbf{X}}|<\infty$ for all $(\xi,t)\in Q$, where $Q\subset\mathbb{R}^{d}\times[0,\infty)$ is a subset with full measure, then $\mathbf{X}$ is referred to as an integrable solution process.
We note that (2.14) can also be viewed as a recursion defined on a Doubly Stochastic Yule Cascade (DSY): $\mathbb{Y}=\{Y_{v}\}_{v\in\mathbb{T}}$ indexed by a binary tree $\mathbb{T}=\{\theta\}\bigcup\left(\cup_{n=1}^{\infty}\{1,2\}^{n}\right)$ rooted at $\theta$, where $Y_{\emptyset}=Y$ is an exponential random variable with intensity $|\xi|^{2}$, $Y_{1}$ and $Y_{2}$ are, conditionally on $Y_{\emptyset}$, the exponential random variables with intensities $|W_{1}|^{2}$ and $|W_{2}|^{2}$, and so on down the tree $\mathbb{T}$, see Section 3 and Section 3 for a more complete definition. In this context, the following two parameters, one stochastic and the other deterministic, play an essential role in the construction of a well-defined solution process.
Definition 2.2.
(i) The (stochastic) explosion time is defined by
$S=S_{\xi}=\underset{n\geq 0}{\mathop{\sup}}\,\underset{|v|=n}{\mathop{\min}}\,\sum\limits_{j=0}^{n}Y_{v|j},$ (2.15)
where for each vertex $v\in\mathbb{T}$, we denote by $|v|$ the genealogical height of vertex $v$, and $v|j$ denotes the truncation up to the $j$’th generation with the convention that $v|0=\theta$. (ii) The critical parameter given by
$t_{c}(\chi_{0})=\sup\{t\geq 0:\mathbb{E}_{\xi}[|{\mathbf{X}}|\mathbbm{1}_{[S>t]}]<\infty\text{~~a.e.~~}\xi\in\mathbb{R}^{d}\}$ (2.16)
is referred to as the integrability parameter for initial data $\chi_{0}$.
The existence of a solution process is assured by the construction of the minimal solution process defined by the solution process constructed in the proof of the following proposition.
Proposition 2.3.
For each measurable function $\chi_{0}:\mathbb{R}^{d}\to\mathbb{C}^{d}$, a solution process exists.
Proof.
Since, for fixed $t<S$ only finitely many branchings can occur, $\mathbf{X}(\xi,t)$ is well-defined by backward iteration of the recursion (2.14). For $t\geq S$ define $\mathbf{X}(\xi,t)=0$. ∎
Corollary 2.4.
If $S=\infty$ almost surely, then the minimal solution process is the unique solution process.
Proof.
The no-explosion event $S=\infty$ a.s. implies that for any $t\geq 0$, the random tree implicit to the iteration of the recursion (2.14) is a.s. finite. In particular, since any solution process can be factored as an ${\odot}_{\xi}$}-product, the minimality follows by induction. ∎
Remark 2.5.
While the primary focus of the present paper is on integrable solution processes, there may be other relationships between the deterministic solutions and the stochastic solution processes, e.g., Cauchy principle value integrals, that justify the analysis of non-integrable solution processes.
In the explosive cases (i.e. $S<\infty$ a.s.), non-minimal solution processes may exist, particularly in the case of scalar mean flow equations with multiplicative non-linearity, [20, 19]. Crucially, in such cases the following notion of hyper-explosion can play an essential role.
Definition 2.6.
The (stochastic) hyper-explosion time is defined by
$L=L_{\xi}=\underset{n\geq 0}{\mathop{\sup}}\,\underset{|v|=n}{\mathop{\max}}\,\sum\limits_{j=0}^{n}Y_{v|j},$ (2.17)
A DSY cascade is called hyper-explosive if $\mathbb{P}[L<\infty]>0$, and the event $L<\infty]$ is called hyper-explosion.
Remark 2.7.
Clearly, hyper-explosive cascades are explosive. Moreover, by the arguments involving the evolution of $\mathbb{P}(L<t)$ and $\mathbb{P}(S>t)$, see e.g. [14], one can show that in the case of DSY described in Section 3, both explosion and hyper-explosion are zero-one events for any initial state $\xi$.
Proposition 2.8.
Let $\chi_{0}:\mathbb{R}^{d}\to\mathbb{C}^{d}$ be a measurable function and let $\mathbf{X}$ be a corresponding integrable solution process. Then the function $\chi(\xi,t)=\mathbb{E}_{\xi}{\mathbf{X}}$ is well-defined and measurable on $Q$ with the integrability property
$\int_{0}^{t}|\xi|^{2\gamma}{{e}^{-|\xi|^{2\gamma}s}}{\int_{{{\mathbb{R}}^{d}}}{|{{\chi}}(\eta,s){{\odot}_{\xi}}{{\chi}}(\xi-\eta,s)|H(\eta|\xi)d\eta ds}<\infty}\ \ \ \forall\,(\xi,t)\in Q.$ (2.18)
Moreover, $\chi$ satisfies (nFNS) everywhere in $Q$.
Proof.
The proof is very similar to [18, Prop. 2.2]. Since $\mathbf{X}(\xi,t)\in L^{1}(\mathbb{P}_{\xi})$, the function $\chi(\xi,t)=\int_{\Omega}\mathbf{X}(\xi,t,\omega)\mathbb{P}_{\xi}(d\omega)$ is measurable on $Q$ (see [24, Thm 2.39]). To show (2.18), we take the expected value of the magnitude of both side of (2.14):
$\infty>{{\mathbb{E}}_{\xi}}|\mathbf{X}|=|{{\chi}_{0}}(\xi)|{{e}^{-|\xi|^{2\gamma}t}}+\int_{0}^{t}{|\xi|^{2\gamma}{{e}^{-|\xi|^{2\gamma}s}}\int_{{{\mathbb{R}}^{d}}}{{{\mathbb{E}}_{\xi}}\left[\left.|\mathbf{X}^{(1)}{{\odot}_{\xi}}\mathbf{X}^{(2)}|\,\right|\,W_{1}=\eta,T_{0}=s\right]H(\eta|\xi)d\eta ds}}$
where $\mathbf{X}^{(k)}=\mathbf{X}^{(k)}(W_{k},t-T)$. Using the conditional independence of $\mathbf{X}^{(1)}$ and $\mathbf{X}^{(2)}$ and the linearity of the expectation, we obtain that for a.e. $(\eta,s)\in\mathbb{R}^{d}\times[0,t)$,
${{{\mathbb{E}}_{\xi}}\left[\left.|\mathbf{X}^{(1)}{{\odot}_{\xi}}\mathbf{X}^{(2)}|\,\right|\,W_{1}=\eta,T=s\right]}\geq\left|{{{\mathbb{E}}_{\xi}}\left[\left.\mathbf{X}^{(1)}{{\odot}_{\xi}}\mathbf{X}^{(2)}\,\right|\,W_{1}=\eta,T=s\right]}\right|$
$=\left|{{\mathbb{E}}_{{\eta}}}{{\mathbf{X}}^{(1)}}(\eta,t-s){\odot}_{\xi}{{\mathbb{E}}_{{\xi-\eta}}}{{\mathbf{X}}^{(2)}}(\xi-\eta,s)\right|=|\chi(\eta,t-s){{\odot}_{\xi}}\chi(\xi-\eta,t-s)|,$
and thus (2.18) follows. To show $\chi=\mathbb{E}_{\xi}\mathbf{X}$ satisfies (nFNS), we take expectation of both sides of (2.14) (with $(\xi,t)\in Q$):
$\chi(\xi,t)={{\mathbb{E}}_{\xi}}[\mathbf{X}(\xi,t)]={{\chi}_{0}}(\xi){{e}^{-|\xi|^{2\gamma}t}}+\int_{0}^{t}{|\xi|^{2\gamma}{{e}^{-|\xi|^{2\gamma}s}}\int_{{{\mathbb{R}}^{d}}}{{{\mathbb{E}}_{\xi}}\left[\left.\mathbf{X}^{(1)}{{\odot}_{\xi}}\mathbf{X}^{(2)}\,\right|\,W_{1}=\eta,T=s\right]H(\eta|\xi)d\eta ds}}$
Thus, (nFNS) holds in $Q$. ∎
Special geometry on the 2-dimensional plane leads to a more specific representation of the $\odot$-product as follows.
Proposition 2.9.
Let $f,g:\mathbb{R}^{2}\to\mathbb{C}^{2}$ be vector fields that satisfy the orthogonality condition:
$\xi\cdot f(\xi)=\xi\cdot g(\xi)=0\ \ \ \forall\xi\in\mathbb{R}^{2}\backslash\{0\}.$
For $\xi,\eta,\zeta\in\mathbb{R}^{2}$, $\xi=\eta+\zeta$, one has
$f(\eta)\odot_{\xi}g(\zeta)=\frac{i}{2}\textup{sign}(\xi\times\eta)(f(\eta)\cdot e_{\eta^{\perp}})(g(\zeta)\cdot e_{\zeta^{\perp}})\sin(\theta_{\xi,\eta}-\theta_{\xi,\zeta})e_{\xi^{\perp}}.$ (2.19)
Here, $a\times b=a^{\perp}\cdot b=a_{1}b_{2}-a_{2}b_{1}$ if $a=(a_{1},a_{2})$ and $b=(b_{1},b_{2})$; $\theta_{u,v}\in[0,\pi]$ is the angle between two vectors $u$ and $v$; $\xi^{\perp}$ denotes the rotation by $90^{\circ}$ of $\xi$, i.e. $\xi^{\perp}=\left[\begin{matrix}0&-1\\ 1&0\\ \end{matrix}\right]\xi$.
Proof.
Let $\tilde{f}(\xi)=f(\xi)\cdot e_{\xi^{\perp}}$ and $\tilde{g}(\xi)=g(\xi)\cdot e_{\xi^{\perp}}$. By the bilinearity of the $\odot$-product,
$f(\eta)\odot_{\xi}g(\zeta)=\tilde{f}(\eta)\tilde{g}(\zeta)e_{\eta^{\perp}}\odot_{\xi}e_{\zeta^{\perp}}.$
By the definition of the $\odot$-product,
$e_{\eta^{\perp}}\odot_{\xi}e_{\zeta^{\perp}}$ $=$ $\frac{1}{2}(e_{\eta^{\perp}}\otimes_{\xi}e_{\zeta^{\perp}}+e_{\zeta^{\perp}}\otimes_{\xi}e_{\eta^{\perp}})$
$=$ $-\frac{i}{2}\left((e_{\xi}\cdot e_{\eta^{\perp}})\pi_{\xi^{\perp}}e_{\zeta^{\perp}}+(e_{\xi}\cdot e_{\zeta^{\perp}})\pi_{\xi^{\perp}}e_{\eta^{\perp}}\right)$
$=$ $-\frac{i}{2}\left(\cos\theta_{\xi,\eta^{\perp}}\cos\theta_{\xi^{\perp},\zeta^{\perp}}+\cos\theta_{\xi,\zeta^{\perp}}\cos\theta_{\xi^{\perp},\eta^{\perp}}\right)e_{\xi^{\perp}}$
$=$ $-\frac{i}{2}\left(\cos\theta_{\xi,\eta^{\perp}}\cos\theta_{\xi,\zeta}+\cos\theta_{\xi,\zeta^{\perp}}\cos\theta_{\xi,\eta}\right)e_{\xi^{\perp}}.$
Figure 1: Two halves of the plane
Based at the origin, the vector $\xi$ partitions the plane into two halves, labeled as and in Figure 1.
① $=$ $\{\eta\in\mathbb{R}^{2}:\ \xi\times\eta=\xi^{\perp}\cdot\eta>0\}$
② $=$ $\{\eta\in\mathbb{R}^{2}:\ \xi\times\eta=\xi^{\perp}\cdot\eta<0\}$
If $\eta\in$① then $\xi\times\eta>0$. We have $\cos\theta_{\xi,\eta^{\perp}}=-\sin\theta_{\xi,\eta}$ and $\cos\theta_{\xi,\zeta^{\perp}}=\sin\theta_{\xi,\zeta}$. Thus
$e_{\eta^{\perp}}\odot_{\xi}e_{\zeta^{\perp}}=-\frac{i}{2}\sin(\theta_{\xi,\zeta}-\theta_{\xi,\eta})e_{\xi^{\perp}}=\frac{i}{2}\sin(\theta_{\xi,\eta}-\theta_{\xi,\zeta})e_{\xi^{\perp}}.$
If $\eta\in$② then $\xi\times\eta<0$. We have $\cos\theta_{\xi,\eta^{\perp}}=\sin\theta_{\xi,\eta}$ and $\cos\theta_{\xi,\zeta^{\perp}}=-\sin\theta_{\xi,\zeta}$. Thus
$e_{\eta^{\perp}}\odot_{\xi}e_{\zeta^{\perp}}=-\frac{i}{2}\sin(\theta_{\xi,\eta}-\theta_{\xi,\zeta})e_{\xi^{\perp}}.$
For $\textup{(NSE)}_{d,\gamma}$ to be non-trivial, it is necessary that $d\geq 2$. Upon removal of the pressure and the divergence-free constraint, one arrives at the fractional Burgers equation, a well-known toy model for $\textup{(NSE)}_{d,\gamma}$, especially in the case $d=1$. Namely,
$u_{t}+(-\Delta)^{\gamma}u+\frac{1}{2}(u^{2})_{x}=0,\ \ u(\cdot,0)=u_{0},$ (2.20)
where $u=u(x,t)$, $x\in\mathbb{R}$, $t\geq 0$. For the corresponding range of values of $\gamma$, namely $\gamma\in(\frac{1}{2},\frac{3}{4})$, one can also interpret the function $\chi(\xi,t)=\frac{c_{0}\hat{u}(\xi,t)}{h(\xi)}$ probabilistically as the expected value of a solution process $\mathbf{X}$ defined in a similar manner as (2.14):
[…]
The fractional Navier-Stokes equations ($d\geq 2$) and the fractional Burgers equation ($d=1$) have essentially the same underlying stochastic structure called a doubly stochastic Yule (DSY) cascade, to be defined in the next section.
3 $(d,\gamma)$-DSY cascade
Let $\mathbb{T}=\{\emptyset\}\bigcup\left(\cup_{n=1}^{\infty}\{1,2\}^{n}\right)$ be the full binary tree rooted at $\emptyset$. The notion of a doubly stochastic Yule (DSY) cascade was introduced in [15] as a tree-indexed family of random variables $\{Y_{v}=\lambda^{-1}_{v}T_{v}\}_{v\in\mathbb{T}}$, where $\{\lambda_{v}\}_{v\in\mathbb{T}}$ is a family of positive random variables independent of the family of i.i.d. mean-one exponentially distributed random variables $\{T_{v}\}_{v\in\mathbb{T}}$.
Definition 3.1.
For $d\geq 1$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$, a $(d,\gamma)$-DSY cascade is a tree-index family $\{Y_{v}=\lambda^{-1}_{v}T_{v}\}_{v\in\mathbb{T}}$ in which $\{T_{v}\}_{v\in\mathbb{T}}$ is an i.i.d. mean-one exponentially distributed family of random variables and $\lambda_{v}=|W_{v}|^{2\gamma}$, where $\{W_{v}\}_{v\in\mathbb{T}}$ is a family of random vectors in $\mathbb{R}^{d}$ distributed as follows:
  1. (i)
    $W_{\emptyset}=\xi,$
  2. (ii)
    Given $W_{v}$, $\lambda_{v}^{-1}T_{v}$ is exponentially distributed with intensity $|W_{v}|^{2\gamma}$, $W_{v1}$ is distributed as $H(\cdot|W_{v})$ and $W_{v2}=W_{v}-W_{v1}$,
  3. (iii)
    Given $W_{v1}$ and $W_{v2}$, the sub-families $\{W_{v1\sigma}\}_{\sigma\in\mathbb{T}}$ and $\{W_{v2\sigma}\}_{\sigma\in\mathbb{T}}$ are independent of each other.
The DSY cascade associated with the fractional Navier-Stokes equations arises when recursively applying formula (2.14). Figure 2 is a cascade figure that illustrates the families $\{{Y}_{v}\}_{v\in\mathbb{T}}$ and $\{W_{v}\}_{v\in\mathbb{T}}$ induced from the recursion (2.14). Although the definition of the $\otimes$-product requires $d\geq 2$, the DSY cascade can be defined for any $d\geq 1$.
Figure 2: Cascade figure that illustrates $\{{Y}_{v}\}_{v\in\mathbb{T}}$ and $\{W_{v}\}_{v\in\mathbb{T}}$.
For each vertex $v\in\mathbb{T}$, we denote by $|v|$ the genealogical height of vertex $v$, and $v|j$ denotes the truncation up to the $j$’th generation with the convention that $v|0=\theta$. Incorporating the time evolution of the $(d,\gamma)$ cascade $\{Y_{v}\}_{v\in\mathbb{T}}$, the recursion (2.14) terminates at a vertex $v\in\mathbb{T}$ if and only if
$\sum_{j=0}^{|v|-1}Y_{v|j}<t\leq\sum_{j=0}^{|v|}Y_{v|j}\,.$
Such a vertex will be referred to as a $t$-leaf. The continuous-parameter Markov process of sets of $t$-leaves is denoted by
$\partial{V}(\xi,t)=\left\{v\in\mathbb{T}:\sum_{j=0}^{|v|-1}Y_{v|j}<t\leq\sum_{j=0}^{|v|}Y_{v|j}\right\}.$ (3.1)
We denote the set of internal vertices by
$\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)=\left\{u\in\mathbb{T}:\ \sum_{j=0}^{|u|}Y_{u|j}<t\right\},$ (3.2)
and the $t$-subtree associated with $\{Y_{v}\}_{v\in\mathbb{T}}$ by
$V(\xi,t)=\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)\cup\partial{V(\xi,t)}.$ (3.3)
Definition 3.2.
Let $S=S_{\xi}$ be the explosion time defined in (2.15). The DSY cascade is said to be non-explosive if $S=\infty$ a.s. Otherwise, it is said to be explosive.
The main purpose of this section is to give criteria on explosion and non-explosion in terms of the dimension $d$ and the fractional power $\gamma$ (Section 3.2).
Remark 3.3.
Whenever $V(\xi,t)$ is a finite set, i.e. on the event $[S>t]$, one has
$|\partial V(\xi,t)|=|\hskip-3.61371pt\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)|+1,\quad|V(\xi,t)|=2|\hskip-3.61371pt\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)|+1~~t\geq 0,$ (3.4)
On the nonexplosion event $[S>t]$, the solution process $\mathbf{X}$ at (2.14) is a $\odot$-product, in a suitable order, of the initial data $\chi_{0}$ evaluated on the $t$-leaves $\partial{V(\xi,t)}$ of the tree that “reached” time $t$. For example, in the tree realization in Figure 3,
$\partial V(\xi,t)=\{11,121,122,21,22\}$
$\mathbf{X}(\xi,t)=(\chi_{0}(W_{11})\odot_{W_{1}}(\chi_{0}(W_{121})\odot_{W_{12}}\chi_{0}(W_{122})))\odot_{\xi}(\chi_{0}(W_{21})\odot_{W_{2}}\chi_{0}(W_{22}))$
Figure 3: Cascade that is nonexplosive by time $t$.
3.1 Distribution of the ratios
In the following proposition, $\overleftarrow{v}$ denotes the parent vertex of $v$.
Proposition 3.4.
[…]
  • For $d=1$,
    $g_{1,\gamma}(r)=c_{1,\gamma}(|r-r^{2}|^{2\gamma-2}+(r+r^{2})^{2\gamma-2})$ (3.5)
    where $c_{1,\gamma}>0$ is the constant given by (2.10).
  • For $d\geq 2$,
    $g_{d,\gamma}(r)={{C}_{d,\gamma}}\frac{r^{2\gamma-2}}{{{\left({{r}^{2}}+1\right)}^{a}}}\ {{}_{2}}{{F}_{1}}\left(\frac{a}{2},\frac{a+1}{2};\frac{d}{2};\frac{4{{r}^{2}}}{{{\left({{r}^{2}}+1\right)}^{2}}}\right)$ (3.6)
    where $a=\frac{d+1}{2}-\gamma>0$ and
    ${{{C}}_{d,\gamma}}=\frac{2\Gamma(2\gamma-1)\Gamma{{(a)}^{2}}}{\Gamma\left(\frac{d}{2}\right)\Gamma\left(2a-\frac{d}{2}\right)\Gamma{{\left(\frac{2\gamma-1}{2}\right)}^{2}}}.$
Moreover, for any $s\in\partial\mathbb{T}$, the sequence $R_{s|k},k\geq 1$, is i.i.d.
Proof.
It is sufficient to consider $v=1$. Let $R=|W_{1}|/|\xi|$ and let $k:[0,\infty)\to\mathbb{R}$ be any bounded measurable function. For $d=1$,
$\mathbb{E}[k(R)]$ $=$ $\int_{-\infty}^{\infty}{k\left(\frac{|\eta|}{|\xi|}\right)H(\eta|\xi)d\eta}={{c}_{1,\gamma}}\int_{-\infty}^{\infty}{k\left(\frac{|\eta|}{|\xi|}\right)\frac{|\eta{{|}^{2\gamma-2}}|\xi-\eta{{|}^{2\gamma-2}}}{|\xi{{|}^{4\gamma-3}}}d\eta}$
$=$ ${{c}_{1,\gamma}}\int_{0}^{\infty}{k\left(\frac{\eta}{|\xi|}\right)\frac{{{\eta}^{2\gamma-2}}}{|\xi{{|}^{4\gamma-3}}}\left(|\xi-\eta{{|}^{2\gamma-2}}+|\xi+\eta{{|}^{2\gamma-2}}\right)d\eta}$
$=$ ${{c}_{1,\gamma}}\int_{0}^{\infty}{k\left(r\right){{r}^{2\gamma-2}}\left(|1-r{{|}^{2\gamma-2}}+(1+r)^{2\gamma-2}\right)dr}.$
This gives the probability density function (3.5). It is easy to check that the inverse ratio $\frac{1}{R}$ has a probability density function
$\tilde{g}_{1,\gamma}(r)={{g}_{1,\gamma}}\left(\frac{1}{r}\right)\frac{1}{{{r}^{2}}}=\frac{{{g}_{1,\gamma}}(r)}{{{r}^{6\gamma-4}}}.$
Thus, $R$ is symmetrically distributed on the multiplicative group $(0,\infty)$ if and only if $\gamma=\frac{2}{3}$. For $d\geq 2$,
$\mathbb{E}[k(R)]=\int_{{{\mathbb{R}}^{d}}\times{{\mathbb{R}}^{d}}}{k\left(\frac{|\eta|}{|\xi|}\right)H(\eta|\xi)}d\eta={{c}_{d,\gamma}}\int_{{{\mathbb{R}}^{d}}}{k\left(\frac{|\eta|}{|\xi|}\right)\frac{|\xi{{|}^{d+2-4\gamma}}}{|\eta{{|}^{d+1-2\gamma}}|\xi-\eta{{|}^{d+1-2\gamma}}}d\eta}$
Let $u=\xi/|\xi|$ and $\tilde{u}=\eta/|\xi|$. Then one can simplify $\mathbb{E}[k(R)]$ as
$\mathbb{E}[k(R)]={{c}_{d,\gamma}}\int_{{{\mathbb{R}}^{d}}}{k(|\tilde{u}|)\frac{1}{|\tilde{u}{{|}^{d+1-2\gamma}}|u-\tilde{u}{{|}^{d+1-2\gamma}}}d\tilde{u}}$. (3.7)
[…]
$\mathbb{E}[k(R)]$ $=$ $\bar{C}_{d,\gamma}\int_{0}^{\infty}{\int_{0}^{\pi}{k(r){{r}^{2\gamma-2}}\frac{\sin^{d-2}\phi}{{{(1-2r\cos\phi+{{r}^{2}})}^{\frac{d+1-2\gamma}{2}}}}d\phi dr}}$
where
${\bar{C}_{d,\gamma}}=|{{\mathbb{S}}^{d-2}}|{{c}_{d,\gamma}}=\frac{2}{\sqrt{\pi}}\frac{\Gamma(2\gamma-1)\Gamma{{\left(a\right)}^{2}}}{\Gamma\left(\frac{d-1}{2}\right)\Gamma\left(2a-\frac{d}{2}\right)\Gamma{{\left(\frac{2\gamma-1}{2}\right)}^{2}}}.$ (3.8)
Then the p.d.f. of $R$ is
$g_{d,\gamma}(r)=\bar{C}_{d,\gamma}{{r}^{2\gamma-2}}\int_{0}^{\pi/2}{{{\sin}^{d-2}}\phi\left({{(1-2r\cos\phi+{{r}^{2}})}^{-a}}+{{(1+2r\cos\phi+{{r}^{2}})}^{-a}}\right)d\phi}$
With the aid of Mathematica (or using [25, Formula 3.682, p.415]) to evaluate the above integral, one obtains
$g_{d,\gamma}(r)$ $=$ $\sqrt{\pi}{\bar{C}_{d,\gamma}}\frac{\Gamma\left(\frac{d-1}{2}\right)}{\Gamma\left(\frac{d}{2}\right)}\frac{r^{2\gamma-2}}{{{\left({{r}^{2}}+1\right)}^{a}}}\ {{}_{2}}{{F}_{1}}\left(\frac{a}{2},\frac{a+1}{2};\frac{d}{2};\frac{4{{r}^{2}}}{{{\left({{r}^{2}}+1\right)}^{2}}}\right)$
$=$ ${C}_{d,\gamma}\frac{r^{2\gamma-2}}{{{\left({{r}^{2}}+1\right)}^{a}}}\ {{}_{2}}{{F}_{1}}\left(\frac{a}{2},\frac{a+1}{2};\frac{d}{2};\frac{4{{r}^{2}}}{{{\left({{r}^{2}}+1\right)}^{2}}}\right)$
The proof that the sequence $\{R_{s|k}\}_{k\geq 1}$ is i.i.d. is a similar calculation by induction, conditioning and using the Markov property along rays. The details are left to the reader. ∎
Definition 3.5.
(see [16, Def. 2.5, Lem. 7.1]) The DSY $\{Y_{v}\}_{v\in\mathbb{T}}$ is said to be self-similar if for any initial state $W_{\emptyset}=\xi$ and for any $s\in\partial\mathbb{T}$, the ratios $\{R_{s|j}=|W_{s|j}|/|W_{s|j-1}|\}_{j\geq 1}$ are i.i.d. with a common distribution independent of $s$.
Corollary 3.6.
The DSY cascade $\{Y_{v}\}_{v\in\mathbb{T}}$ is self-similar. Moreover, the ratio $R_{v}=\frac{|W_{v}|}{|W_{\overleftarrow{v}}|}$ is symmetrically distributed on the multiplicative group $(0,\infty)$, i.e. having the same distribution as $R_{v}^{-1}$, if and only if $\gamma=\frac{d+3}{6}$.
Proof.
By Section 3.1, the ratios $\{R_{s|k}\}_{k\geq 1}$ are i.i.d. Thus, $\{Y_{v}\}_{v\in\mathbb{T}}$ is self-similar. The inverse ratio $\frac{1}{R}$ has a probability density function
$\tilde{g}_{d,\gamma}(r)={{g}_{d,\gamma}}\left(\frac{1}{r}\right)\frac{1}{{{r}^{2}}}=\frac{{{g}_{d,\gamma}}(r)}{{{r}^{4\gamma-2a-2}}}=\frac{g_{d,\gamma}(r)}{r^{6\gamma-d-3}}.$
Therefore, $R$ is symmetrically distributed on the multiplicative group $(0,\infty)$ if and only if $\gamma=\frac{d+3}{6}$. ∎
Proposition 3.7.
Let $d\geq 2$. For each $v\in\mathbb{T}$, let $\Phi_{v1}\in[0,\pi]$ be the angle between $W_{v}$ and $W_{v1}$, and $\Phi_{v2}\in[0,\pi]$ be the angle between $W_{v}$ and $W_{v2}$. The joint distribution density of the random pair $(\Phi_{1},\Phi_{2})\in\triangle:=\{(\phi_{1},\phi_{2}):\ \phi_{1},\phi_{2}\geq 0,\ \phi_{1}+\phi_{2}\leq\pi\}$ conditionally given $W_{v}$ is given by
$f_{d,\gamma}(\phi_{1},\phi_{2})=\bar{C}_{d,\gamma}(\sin\phi_{1}\sin\phi_{2})^{2\gamma-2}{\sin}^{d+1-4\gamma}(\phi_{1}+\phi_{2})$ (3.9)
where the constant $\bar{C}_{d,\gamma}$ is given by (3.8).
Proof.
It is sufficient to consider the root vertex $v=\theta$. Let $W_{0}=\xi$, $W_{1}=\eta$, $\Phi_{1}=\phi_{1}$, $\Phi_{2}=\phi_{2}$, $u=\xi/|\xi|$, $\tilde{u}=\eta/|\xi|$. By the law of sines in the triangle made by $u$, $\tilde{u}$, $u-\tilde{u}$, we get
[…]
Thus, the Jacobian of the change of variables from $(r,\phi,w)$ to $(\phi_{1},\phi_{2},w)$ is $\frac{\partial(r,\phi,w)}{\partial(\phi_{1},\phi_{2},w)}=\frac{-\sin\phi_{1}}{\sin^{2}(\phi_{1}+\phi_{2})}$. We obtain
$d\tilde{u}=r^{d-1}\sin^{d-2}\phi\,drd\phi dw=\frac{(\sin\phi_{1}\sin\phi_{2})^{d-1}}{\sin^{d+1}(\phi_{1}+\phi_{2})}d\phi_{1}d\phi_{2}dw.$ (3.12)
Let $k:\triangle\to\mathbb{R}$ be a continuous and bounded function. By (3.7) and (3.12),
$\mathbb{E}[k(\Phi_{1},\Phi_{2})]$ $=$ ${{c}_{d,\gamma}}\int_{{{\mathbb{R}}^{d}}}{k(\phi_{1},\phi_{2})\frac{1}{|\tilde{u}{{|}^{d+1-2\gamma}}|u-\tilde{u}{{|}^{d+1-2\gamma}}}d\tilde{u}}$
$=$ ${{c}_{d,\gamma}}\int_{\mathbb{S}_{u}^{d-2}}{\int_{\triangle}{k({{\phi}_{1}},{{\phi}_{2}}){{(\sin{{\phi}_{1}}\sin{{\phi}_{2}})}^{2\gamma-2}}{{\sin}^{d+1-4\gamma}}({{\phi}_{1}}+{{\phi}_{2}})d{{\phi}_{1}}d{{\phi}_{2}}dw}}$
$=$ ${{{\bar{C}}}_{d,\gamma}}\int_{\triangle}{k({{\phi}_{1}},{{\phi}_{2}}){{(\sin{{\phi}_{1}}\sin{{\phi}_{2}})}^{2\gamma-2}}{{\sin}^{d+1-4\gamma}}({{\phi}_{1}}+{{\phi}_{2}})d{{\phi}_{1}}d{{\phi}_{2}}}.$
Hence, the joint distribution of $(\Phi_{2},\Phi_{2})$ is given by (3.9). ∎
Remark 3.8.
In the classical case $d=3$ and $\gamma=1$, Le Jan and Sznitman [28, Prop. 1.2] remark that the angular vector $(\Phi_{1},\Phi_{2})$ is uniformly distributed in the triangle $\triangle$. It is easy to see that this is the case if and only if $d=3$ and $\gamma=1$.
Remark 3.9.
The symmetry with respect the multiplicative group $(0,\infty)$ of the ratio $R$ noted in Section 3.1 is a consequence of the invariance of the joint distribution of $(R_{1},R_{2})$ with respect to the Kelvin inversion with respect to the unit sphere. This can proved using the joint distribution of the angles given in Section 3.1 as follows. Note that in terms of the angles the Kelvin inversion is the transformation (see Figure 4)
$(\phi_{1},\phi_{2})\rightarrow(\phi_{1},\phi_{3}),\qquad\phi_{3}=\pi-(\phi_{1}+\phi_{2}).$ (3.13)
Since $\sin(\phi_{1}+\phi_{2})=\sin(\phi_{3})$, the joint density of the angles is invariant iff
$d+1-4\gamma=2\gamma-2$
Figure 4: Kelvin inversion $\eta\mapsto\eta^{*}$, where $W_{1}=\eta$ and $\eta^{*}=\frac{1}{|\eta|^{2}}\eta$.
i.e. iff $\gamma=\tfrac{d+3}{6}$. The coordinate transformation (3.13) will be used in the proof of Section 3.2, highlighting the importance of this symmetry in the explosion properties of the DSY cascade $\mathbb{Y}$.
Remark 3.10.
Alternatively, instead of angles $\Phi_{1}$ and $\Phi_{2}$, one can describe the basic triangle formed by $W_{v}$, $W_{v1}$, and $W_{v2}$ in terms of (normalized) side lengths: $1=|W_{v}|/|W_{v}|$, $R_{1}=|W_{v1}|/|W_{v}|$, and $R_{2}=|W_{v2}|/|W_{v}|$, with the joint distribution of $R_{1}$ and $R_{2}$ given by the pdf
$h_{d,\gamma}(r_{1},r_{2})=\tilde{C}_{d,\gamma}\frac{A(1,r_{1},r_{2})^{d-3}}{(r_{1}r_{2})^{d-2\gamma}}{\mathds{1}_{D}}(r_{1},r_{2}),$ (3.14)
where, using Heron’s triangle formula for area,
$A(1,r_{1},r_{2})=\frac{1}{4}\sqrt{(1+r_{1}+r_{2})(1+r_{1}-r_{2})(1+r_{2}-r_{1})(r_{1}+r_{2}-1)}$
is the area of the triangle with sides $1$, $r_{1}$, $r_{2}$. Here,
$D=\{(r_{1},r_{2})\in\mathbb{R}^{2}:r_{1},r_{2}>0,r_{1}+r_{2}>1,|r_{1}-r_{2}|<1\}.$
As we will see, moments of random variables $R_{v}$ will play a defining role in the explosive/non-explosive nature of the corresponding DSY cascades.
Proposition 3.11.
Let $d\geq 2$, $\gamma\in(\frac{1}{2},\frac{d+2}{4})$ and $\bar{C}_{d,\gamma}$ be the constant given by (3.8). One has:
  1. (i)
    For all $b\in\mathbb{R}$, $\mathbb{E}[R^{b}]=\mathbb{E}[R^{\bar{b}}]$ where $\bar{b}=d+3-6\gamma-b$.
  2. (ii)
    Let $b^{*}=\frac{d+3}{2}-3\gamma$. Then
    $\min\limits_{b\in\mathbb{R}}\mathbb{E}[R^{b}]=\mathbb{E}[R^{b^{*}}]={\bar{C}_{d,\gamma}}\int_{0}^{\pi}\int_{0}^{\pi-\phi_{1}}(\sin\phi_{2}\,\sin(\phi_{1}+\phi_{2}))^{\frac{d-1}{2}-\gamma}\sin^{2\gamma-2}\phi_{1}d\phi_{2}d\phi_{1}.$
  3. (iii)
    $\mathbb{E}[R^{b}]\geq 1$ for all $(b,\gamma)\in(0,\infty)\times[\frac{d+3}{6},\frac{d+2}{4})\bigcup\,(-\infty,0)\times(\frac{1}{2},\frac{d+3}{6}]$.
Proof.
(i) Let $a=\frac{d+1}{2}-\gamma>0$ and
$\tilde{F}_{d,\gamma}(r)=\,_{2}F_{1}\left(\frac{a}{2},\frac{a+1}{2},\frac{d}{2},\frac{4r^{2}}{(r^{2}+1)^{2}}\right).$
For any $b\in\mathbb{R}$,
$\mathbb{E}[R^{b}]=\int_{0}^{\infty}C_{d,\gamma}\frac{r^{2\gamma-2+b}}{(r^{2}+1)^{a}}\tilde{F}_{d,\gamma}(r)dr.$ (3.15)
Note that $\tilde{F}_{d,\gamma}(r)=\tilde{F}_{d,\gamma}(1/r)$. Using the change of variable $r\mapsto 1/r$ for the integral at (3.15), we get
$\mathbb{E}[R^{b}]=\int_{0}^{\infty}C_{d,\gamma}\frac{r^{2a-2\gamma-b}}{(r^{2}+1)^{a}}\tilde{F}_{d,\gamma}(r)dr=\mathbb{E}[R^{\bar{b}}].$
(ii) By splitting $\int_{0}^{\infty}=\int_{0}^{1}+\int_{1}^{\infty}$ and using the change of variable $r\mapsto 1/r$ for the last integral, we get
$\mathbb{E}[{{R}^{b}}]=\int_{0}^{1}{\frac{{{C}_{d,\gamma}}}{{{({{r}^{2}}+1)}^{a}}}\left({{r}^{2\gamma-2+b}}+{{r}^{2a-2\gamma-b}}\right){{{\tilde{F}}}_{d,\gamma}}(r)dr}$
By the Cauchy-Schwarz inequality, or since the geometric mean cannot exceed the arithmetic mean, $r^{2\gamma-2+b}+r^{2a-2\gamma-b}\geq 2r^{a-1}$. Thus,
$\mathbb{E}[{{R}^{b}}]\geq\int_{0}^{1}{\frac{{{C}_{d,\gamma}}}{{{({{r}^{2}}+1)}^{a}}}2{{r}^{a-1}}{{{\tilde{F}}}_{d,\gamma}}(r)dr}=\mathbb{E}[{{R}^{b^{*}}}].$ (3.16)
By (3.10), one can rewrite $\mathbb{E}[R^{b^{*}}]$ in terms of the angles as
$\mathbb{E}[R^{b^{*}}]={\bar{C}_{d,\gamma}}\int_{0}^{\pi}\int_{0}^{\pi-\phi_{1}}(\sin\phi_{2}\,\sin(\phi_{1}+\phi_{2}))^{\frac{d-1}{2}-\gamma}\sin^{2\gamma-2}\phi_{1}d\phi_{2}d\phi_{1}.$
(iii) In the case $(b,\gamma)\in(0,\infty)\times[\frac{d+3}{6},\frac{d+2}{4})\bigcup\,(-\infty,0)\times(\frac{1}{2},\frac{d+3}{6}]$, note that $b$ and $\bar{b}$ have different signs. There exists $p\in(1,\infty)$ such that $\frac{b}{p}+\frac{\bar{b}}{p^{\prime}}=0$ where $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. The proof proceeds by noticing the logarithmic convexity of the moment function $x\mapsto\mathbb{E}[R^{x}]$ as follows. By Holder’s inequality,
$\mathbb{E}[{{R}^{b}}]=\mathbb{E}{{[{{R}^{b}}]}^{1/p}}\mathbb{E}{{[{{R}^{{\bar{b}}}}]}^{1/p^{\prime}}}\geq\mathbb{E}[{{R}^{b/p+\bar{b}/p^{\prime}}}]=\mathbb{E}[{{R}^{0}}]=1.$
3.2 Explosion and hyper-explosion of $(d,\gamma)$-DSY cascades
We start by recalling the explosion and non-explosion criteria for self-similar $(d,\gamma)$-DSY cascades from [16]. Note that part (b) in the theorem below is in fact an improvement of [16, Prop. 2.7], which can be applied to general self-similar DSY considered there. We also establish two new conditions for hyperexplosion and nonhyperexplosion, which for simplicity are formulated in the context of the $(d,\gamma)$-DSY, but transfer in a straightforward way to general self-similar DSY considered in [16].
Theorem 3.12.
A self-similar $(d,\gamma)$-DSY cascade is
  1. (a)
    non-explosive if $\mathbb{E}[R^{b}]<1/2$ for some $b>0$;
  2. (b)
    explosive if $\mathbb{E}[\ln R_{\max}]>0$, where $R_{\max}=\max\{R_{1},R_{2}\}$, $R_{1}=\frac{|W_{1}|}{|W_{\theta}|}$, $R_{2}=\frac{|W_{2}|}{|W_{\theta}|}$.
  3. (c)
    If $\mathbb{E}[\ln R_{\min}]<0$, where $R_{\min}=\min\{R_{1},R_{2}\}$ then the cascade is a.s. not hyperexplosive.
  4. (d)
    If $\mathbb{E}[R^{-2\gamma\alpha}]<\frac{1}{2}$ for some $\alpha\in(0,1]$ then the DSY cascade is a.s. hyperexplosive.
Proof.
Part (a) is [16, Prop. 2.6]. For part (b) the proof is similar to that of [16, Prop. 2.7]. Namely, consider the ”greedy” random path (ray of the DSY tree) where at each generation $n$, we select a sibling with largest ratio $R_{v}$, i.e, inductively, at vertex $v$, select a next vertex $v1$ if $R_{v1}>R_{v2}$ and vertex $v2$ otherwise. In this way, the DSY tree contains a greedy ray $\bar{s}=(s_{1},s_{2},\dots)\in\{1,2\}^{\mathbb{N}}=\partial\mathbb{T}$ with associated ”time”
$\tau_{\bar{s}}=\sum_{j=0}^{\infty}\frac{T_{\bar{s}|j}}{|W_{\bar{s}|j}|^{2\gamma}}=\sum_{j=0}^{\infty}\frac{T_{\bar{s}|j}}{\prod_{k=0}^{j}R_{\max,k}^{2\gamma}}\geq S,$
where $R_{\max,j+1}=\max\{R_{[\bar{s}|j]1},R_{[\bar{s}|j]2}\}$ are the iid copies of $R_{\max}$. If $\mathbb{E}[\ln(R_{\max})]=\mu<0$, then by the law of large numbers,
$\frac{\ln(\prod_{k=0}^{j}R_{\max,k})}{j}=\frac{\sum_{k=0}^{j}\ln(R_{\max,k})}{j}\to\mu,\qquad\mbox{a.s. as}\ j\to\infty,$
and therefore, a.s. there will exist a (random) $J\in\mathbb{N}$ such that for $j>J$
$\frac{\sum_{k=0}^{j}\ln(R_{\max,k})}{j}>\frac{\mu}{2},$
and therefore, for $j\geq J$,
$\frac{1}{\prod_{k=0}^{j}R_{\max,k}^{2\gamma}}<\frac{1}{\alpha^{j}},\qquad\alpha=e^{\mu\gamma}>1.$
Consequently, since $\{T_{\bar{s}|j}\}_{j\in\mathbb{N}}$ are iid exponential random variables with intensity 1 and the series $\sum_{j\geq J}1/\alpha^{j}<\infty$, we have
$\tau_{\bar{s}}\leq\sum_{j=0}^{J}\frac{T_{\bar{s}|j}}{\prod_{k=0}^{j}R_{\max,k}^{2\gamma}}+\sum_{j=J+1}^{\infty}\frac{T_{\bar{s}|j}}{\alpha^{j}}\leq\sum_{j=0}^{J}\frac{T_{\bar{s}|j}}{\prod_{k=0}^{j}R_{\max,k}^{2\gamma}}+\sum_{j=1}^{\infty}\frac{T_{\bar{s}|j}}{\alpha^{j}}<\infty\quad a.s.,$
since the first term is a finite sum and the second has finite expected value. Thus, $S<\infty$ a.s.
Part (c) is proven similarly to Part (b) except this time one follows the ray $\underline{s}$ of minimal siblings instead of maximal, and, by the law of large numbers, the corresponding time will be bounded below by the tail of the a.s. divergent series $\sum_{j>J}{T_{\underline{s}|j}}\alpha^{j}$, where $\alpha=e^{\mu/2}$ with $\mu=-\mathbb{E}[\ln R_{\min}]>0$.
For part (d), due to the self-similarity of the DSY cascade, one can assume $|\xi|=1$. Let
$L_{n}=\max\limits_{|v|=n}\sum\limits_{j=0}^{n}\frac{T_{v|j}}{|W_{v|j}|^{2\gamma}}$
be the maximum length of all paths of length $n$. One has
$L=\lim\limits_{n\to\infty}L_{n}$
and
$L_{n+1}=T_{0}+\max\{L^{(1)}_{n}R_{1}^{-2\gamma},\,L^{(2)}_{n}R_{2}^{-2\gamma}\}.$ (3.17)
Equation (3.17) implies
$L_{n+1}\leq T_{0}+L^{(1)}_{n}R_{1}^{-2\gamma}+L^{(2)}_{n}R_{2}^{-2\gamma}.$ (3.18)
In view of the inequality
$(a+b)^{\alpha}\leq a^{\alpha}+b^{\alpha}\ \ \ \forall\,a,b>0,\alpha\in(0,1],$
it follows from (3.18) that for any $\alpha\in(0,1]$,
$L^{\alpha}_{n+1}\leq T^{\alpha}_{0}+(L^{(1)}_{n})^{\alpha}R_{1}^{-2\gamma\alpha}+(L^{(2)}_{n})^{\alpha}R_{2}^{-2\gamma\alpha}.$ (3.19)
Take the expectation of both sides:
$\mathbb{E}[L^{\alpha}_{n+1}]$ $\leq$ $\mathbb{E}[T^{\alpha}_{0}]+\mathbb{E}[(L^{(1)}_{n})^{\alpha}R_{1}^{-2\gamma\alpha}]+\mathbb{E}[(L^{(2)}_{n})^{\alpha}R_{2}^{-2\gamma\alpha}].$
By the self-similarity of the cascade, the vector $(L_{n}^{(1)},L_{n}^{(2)})$ is independent of the vector $(R_{1},R_{2})$. Therefore, the above inequality can be written as
$\mathbb{E}[L^{\alpha}_{n+1}]$ $\leq$ $A+2\mathbb{E}[R_{1}^{-2\gamma\alpha}]\,\mathbb{E}[L^{\alpha}_{n}].$ (3.20)
By the estimate (3.20), the sequence $\{\mathbb{E}[L_{n}^{\alpha}]\}_{n\geq 0}$ is bounded. Consequently,
$\mathbb{E}[L^{\alpha}]=\lim\mathbb{E}[L_{n}^{\alpha}]<\infty.$
where $L=\lim L_{n}$ is the longest random path on the DSY cascade. Therefore, $L<\infty$ almost surely. ∎
Remark 3.13.
Section 3.2, in particular, parts (b)-(c) can be straightforwardly adapted to the general case of self-similar DSY cascades as defined in [16]. In the case of fractional Navier-Stokes equations, $d\geq 2$, Section 3.1 provides explicit expressions for the expectations involved in Section 3.2. Indeed, given that $\Phi_{1}$ and $\Phi_{2}$ are angles in the triangle with sides of lengths $1$, $R_{1}$, $R_{2}$ (with $\Phi_{1}$ opposite to $R_{2}$ and $\Phi_{2}$ opposite to $R_{1}$) one can use the law of sines to write $R_{1}=r_{1}(\Phi_{1},\Phi_{2})=\sin{\Phi_{2}}/\sin(\Phi_{1}+\Phi_{2})$ and $R_{2}=r_{2}(\Phi_{1},\Phi_{2})=\sin{\Phi_{1}}/\sin(\Phi_{1}+\Phi_{2})$, obtaining
$\mathbb{E}[R_{1}^{-2\gamma\alpha}]$ $=$ $\iint_{\triangle}r_{1}(\phi_{1},\phi_{2})^{-2\gamma\alpha}f_{d,\gamma}(\phi_{1},\phi_{2})\,d\phi_{2}d\phi_{1}$
$=$ $\bar{C}_{d,\gamma}\int\limits_{0}^{\pi}\int\limits_{0}^{\pi-\phi_{1}}(\sin\phi_{1})^{2\gamma-2}(\sin\phi_{2})^{2\gamma(1-\alpha)}(\sin(\phi_{1}+\phi_{2}))^{d+1-4\gamma+2\gamma\alpha}d\phi_{2}d\phi_{1}.$
Moreover, noting that $R_{1}\geq R_{2}$ on the subregion
$(\Phi_{1},\Phi_{2})\in\tilde{\Delta}=\{(\phi_{1},\phi_{2})\in\Delta:\ \phi_{1}\leq\phi_{2},\ \phi_{1},\phi_{2}\geq 0,\ \phi_{1}+\phi_{2}\leq\pi\},$
we have
$\mathbb{E}[\ln R_{\max}]$ $=$ $2\iint\limits_{\tilde{\Delta}}\ln\left(r_{1}(\phi_{1},\phi_{2})\right)f_{d,\gamma}(\phi_{1},\phi_{2})\,d\phi_{2}d\phi_{1}$
$=$ $2\bar{C}_{d,\gamma}\int\limits_{0}^{\pi/2}\int\limits_{\phi_{1}}^{\pi-\phi_{1}}\ln\left(\frac{\sin\phi_{2}}{\sin(\phi_{1}+\phi_{2})}\right)(\sin\phi_{1}\sin\phi_{2})^{2\gamma-2}{\sin}^{d+1-4\gamma}(\phi_{1}+\phi_{2})\,d\phi_{1}d\phi_{2},$
$\mathbb{E}[\ln R_{\min}]$ $=$ $2\iint\limits_{\tilde{\Delta}}\ln\left(r_{2}(\phi_{1},\phi_{2})\right)f_{d,\gamma}(\phi_{1},\phi_{2})\,d\phi_{2}d\phi_{1}$
$=$ $2\bar{C}_{d,\gamma}\int\limits_{0}^{\pi/2}\int\limits_{\phi_{1}}^{\pi-\phi_{1}}\ln\left(\frac{\sin\phi_{1}}{\sin(\phi_{1}+\phi_{2})}\right)(\sin\phi_{1}\sin\phi_{2})^{2\gamma-2}{\sin}^{d+1-4\gamma}(\phi_{1}+\phi_{2})\,d\phi_{1}d\phi_{2},$
Note that since $\gamma\in\left(\frac{1}{2},\frac{d+2}{4}\right)$, the integral in (3.13) is finite if and only if $\alpha\in\left(0,1-\frac{1}{2\gamma}\right)$.
In the case of fractional Navier-Stokes equations, Section 3.2 above leads to the following result.
Theorem 3.14.
Let $d\geq 1$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$. Let $\mathbf{Y}=\{Y_{v}=|W_{v}|^{-2\gamma}T_{v}\}_{v\in\mathbb{T}}$ be the DSY cascade.
  • (a)
    $\mathbf{Y}$ is nonexplosive if $d\geq 2$, $\gamma\in(\frac{1}{2},\frac{d+3}{6})$ and $\mathbb{E}[R^{b^{*}}]<1/2$, where $b^{*}=\frac{d+3}{2}-3\gamma$. Moreover, $b^{*}$ corresponds to the best choice of $b$ in Section 3.2 (a) in the sense that $\mathbb{E}[R^{b^{*}}]=\min_{b>0}\mathbb{E}[R^{b}]$.
  • (b)
    $\mathbf{Y}$ is explosive if $d=1,\gamma\in(\frac{1}{2},\frac{3}{4})$ or $d\geq 2,\gamma\in[\frac{d+3}{6},\frac{d+2}{4})$.
  • (c)
    $\mathbf{Y}$ is non-hyperexplosive if $d\geq 2,\gamma\in(\frac{1}{2},\frac{d+3}{6}]$.
  • (d)
    $\mathbf{Y}$ is hyperexplosive if $d\geq 2$, $\gamma\in(\frac{d+3}{6},\frac{d+2}{4})$ and $\mathbb{E}[R^{b^{*}}]<1/2$. Moreover, $b^{*}$ corresponds to the best choice of $\alpha$ in Section 3.2 (c) in the sense that $\mathbb{E}[R^{b^{*}}]=\min_{\alpha\in(0,1]}\mathbb{E}[R^{-2\gamma\alpha}]$.
Remark 3.15.
The simple ranges of $\gamma$ in Part (b) and (c) are smaller than the ranges of $\gamma$ determined by the general hyperexplosion criterion Section 3.2 (b), (c) (see Figure 6).
Proof of Section 3.2.
Part (a) follows directly from Section 3.2 (a). The optimality of $b^{*}$ follows from Section 3.1 (b). For Part (d), note that $b^{*}=-2\gamma\alpha$ with $\alpha=\frac{3}{2}-\frac{d+3}{4\gamma}\in(0,1]$. By Section 3.2 (d), the DSY cascade is a.s. hyper-explosive.
We will now prove Part (b). In the case $d=1$, we can establish explosion by showing $\mathbb{E}(R_{\max}^{-2\gamma})<1$ which yields explosion by [16, Prop. 2.7]. Let $g_{R_{\max}}$ be the pdf of $R_{\max}$, which is given in Section 3.2. Because $H(\cdot|1)$ is a probability density symmetric with respect to $1/2$, we have
$\mathbb{E}[R_{\max}^{-2\gamma}]=\int\limits_{0}^{\infty}{{{r}^{-2\gamma}}g_{R_{\max}}(r)dr}=\frac{\int\limits_{0}^{\infty}{{{r}^{-2\gamma}}g_{R_{\max}}(r)dr}}{\int\limits_{-\infty}^{\infty}{H(r|1)}dr}=\frac{\int\limits_{1/2}^{\infty}{{{r}^{-2}}|1-r{{|}^{2\gamma-2}}dr}}{\int\limits_{1/2}^{\infty}{|r-{{r}^{2}}{{|}^{2\gamma-2}}dr}}.$
We want to show that
$\int\limits_{1/2}^{\infty}{{{r}^{-2}}|1-r{{|}^{2\gamma-2}}dr}<\int\limits_{1/2}^{\infty}{|r-{{r}^{2}}{{|}^{2\gamma-2}}dr}$
Using the changing of variable $r\mapsto 1/r$ on both sides, we can rewrite the inequality to be shown as
$\int\limits_{0}^{2}{{{r}^{2-2\gamma}}|1-r{{|}^{2\gamma-2}}dr}<\int\limits_{0}^{2}{{{r}^{2-4\gamma}}|1-r{{|}^{2\gamma-2}}dr}.$ (3.24)
The integrals on both sides are then split into $\int_{0}^{1}+\int_{1}^{2}$. Using the change of variable $r\mapsto 2-r$ for the integral $\int_{1}^{2}$, one obtains an inequality equivalent to (3.24)
$\int\limits_{0}^{1}{[{{r}^{a}}+{{(2-r)}^{a}}]{{(1-r)}^{-a}}dr}<\int\limits_{0}^{1}{[{{r}^{b}}+{{(2-r)}^{b}}]{{(1-r)}^{-a}}dr}$ (3.25)
where $a=2-2\gamma\in(1/2,1)$ and $b=2-4\gamma\in(-1,0)$. To show (3.25), it suffices to show that
$r^{a}+(2-r)^{a}<r^{b}+(2-r)^{b}\ \ \ \forall r\in(0,1).$ (3.26)
Because $a\in(1/2,1)$, the function $f(x)=x^{a}$ is a concave function on $(0,\infty)$. Thus,
[…]
Because $b\in(-1,0)$, the function $g(x)=x^{b}$ is a convex function on $(0,\infty)$. Thus,
[…]
Therefore, (3.26) has been proven, and explosion in the case $d=1$ follows.
In the case $d\geq 2$ use Section 3.1 to estimate $\mathbb{E}[\ln(R_{\max})]$ by exploiting the symmetry about $\alpha=(d+3)/6$, one manifestation of which was seen in Section 3.1.
Indeed, given that $\Phi_{1}$ and $\Phi_{2}$ are angles in the triangle with sides of lengths $1$, $R_{1}$, $R_{2}$ (with $\Phi_{1}$ opposite to $R_{2}$ and $\Phi_{2}$ opposite to $R_{1}$), we note that $R_{1}>R_{2}$ when $\Phi_{1}<\Phi_{2}$. Thus, by the absolute continuity of the joint distribution of $(\Phi_{1},\Phi_{2})$ with respect to the Lebesgue measure on $\Delta\{(\phi_{1},\phi_{2}):\ \phi_{1},\phi_{2}\geq 0,\ \phi_{1}+\phi_{2}\leq\pi\}$, we conclude that
$\mathbb{E}[\ln R_{\max}]>\mathbb{E}[\ln R_{1}]>\mathbb{E}[\ln R_{\min}]$ (3.27)
$\mathbb{E}[\ln R_{1}]=2\iint\limits_{\Delta}\ln(r_{1}(\phi_{1},\phi_{1}))f_{d,\gamma}(\phi_{1},\phi_{2})d\phi_{1}d\phi_{2},$
where $R_{1}=r_{1}(\phi_{1},\phi_{2})$ when $\Phi_{1}=\phi_{1}$ and $\Phi_{2}=\phi_{2}$.
By the law of sines, $r_{1}(\phi_{1},\phi_{2})=\sin\phi_{2}/\sin\phi_{3}=\sin\phi_{2}/\sin(\phi_{1}+\phi_{2})$, where $\phi_{3}=\pi-\phi_{1}-\phi_{2}$ is the third angle in the triangle with sides $1$, $r_{1}$, $r_{2}$. Thus, using (3.9):
$\mathbb{E}[\ln R_{1}]=\bar{C}_{d,\gamma}\iint\limits_{\Delta}\ln\left(\frac{\sin\phi_{2}}{\sin(\phi_{1}+\phi_{2})}\right)(\sin\phi_{1}\sin\phi_{2})^{2\gamma-2}{\sin}^{d+1-4\gamma}(\phi_{1}+\phi_{2})\,d\phi_{1}d\phi_{2}\;.$
Since we are interested in the sign of $\mathbb{E}[\ln(R_{\max})]$, we can ignore the normalizing constants and consider
$I=\int\limits_{0}^{\pi}\int\limits_{0}^{\pi-\phi_{1}}\ln\left(\frac{\sin\phi_{2}}{\sin\phi_{3}}\right)(\sin\phi_{1}\sin\phi_{2})^{2\gamma-2}{\sin}^{d+1-4\gamma}\phi_{3}\,d\phi_{2}d\phi_{1}\;,$
with $\phi_{3}=\pi-(\phi_{1}+\phi_{2})$.
Note that $\sin\phi_{2}\geq\sin\phi_{3}$ if either (i) $\frac{\pi}{2}\geq\phi_{2}\geq\phi_{3}(\geq 0)$ or, equivalently, $\phi_{1}\geq\pi-2\phi_{2}$ and $0\leq\phi_{2}\leq\frac{\pi}{2}$ or (ii) if $\phi_{2}>\frac{\pi}{2}$. Thus, denote by
$\Delta_{+}=\{(\phi_{1},\phi_{2})\in\Delta:\ \sin\phi_{2}>\sin\phi_{3}\}=\{(\phi_{1},\phi_{2}):\ \phi_{1},\phi_{2}\geq 0,\ \phi_{1}+\phi_{2}\leq\pi,\ \phi_{1}+2\phi_{2}\geq\pi\}$
and
$\Delta_{-}=\Delta\setminus\Delta_{+}=\{(\phi_{1},\phi_{2}):\ \phi_{2}\geq\phi_{1}\geq 0,\ \phi_{1}<\pi-2\phi_{2}\}$
are the regions where the log inside the integral $I$ is non-negative and negative respectively, so that
$I=I_{+}+I_{-}=\iint_{\Delta_{+}}\dots+\iint_{\Delta_{-}}\dots.$
We can write
$I_{-}=\int\limits_{0}^{\pi}\int\limits_{0}^{\frac{\pi-\phi_{1}}{2}}\ln\left(\frac{\sin\phi_{2}}{\sin\phi_{3}}\right)(\sin\phi_{1}\sin\phi_{2})^{2\gamma-2}{\sin}^{d+1-4\gamma}\phi_{3}\,d\phi_{2}d\phi_{1}\;.$
Change variables in the inner integral: $\theta_{2}=\pi-(\phi_{1}+\phi_{2})=\phi_{3}$, i.e. $\phi_{2}=\pi-(\phi_{1}-\theta_{2})=\theta_{3}$. We obtain
$I_{-}$ $=\int\limits_{0}^{\pi}\int\limits_{\frac{\pi-\phi_{1}}{2}}^{\pi-2\phi_{1}}\ln\left(\frac{\sin\theta_{3}}{\sin\theta_{2}}\right)(\sin\phi_{1}\sin\theta_{3})^{2\gamma-2}{\sin}^{d+1-4\gamma}\theta_{2}\,d\theta_{2}d\phi_{1}$ (3.28)
$=-\int\limits_{0}^{\pi}\int\limits_{\pi-2\phi_{1}}^{\frac{\pi-\phi_{1}}{2}}\ln\left(\frac{\sin\theta_{2}}{\sin\theta_{3}}\right)(\sin\phi_{1}\sin\theta_{2})^{2\gamma-2}{\sin}^{d+1-4\gamma}\theta_{3}\left(\frac{\sin\theta_{2}}{\sin\theta_{3}}\right)^{d+1-4\gamma-(2\gamma-2)}d\theta_{2}d\phi_{1}$
$=-\iint\limits_{\Delta_{+}}\ln\left(\frac{\sin\phi_{2}}{\sin\phi_{3}}\right)(\sin\phi_{1}\sin\phi_{2})^{2\gamma-2}{\sin}^{d+1-4\gamma}\phi_{3}\left(\frac{\sin\phi_{2}}{\sin\phi_{3}}\right)^{d+3-6\gamma}d\theta_{2}d\phi_{1}\;.$
Since on $\Delta_{+}$ we have $\sin\phi_{2}\geq\sin\phi_{3}$, we conclude that when $d+3-6\gamma\leq 0$, i.e. when $\gamma\geq(d+3)/6$, we have:
$I_{-}\geq-I_{+}\quad\mbox{i.e.}\quad I=I_{+}+I_{-}\geq 0\;,$
i.e.
$\mathbb{E}[R_{1}]>0\quad\mbox{iff}\quad\gamma\geq(d+3)/6\;.$
Thus, when $\gamma\geq(d+3)/6$, by (3.27), $\mathbb{E}[\ln R_{\max}]>0$, which, according to Section 3.2 implies explosion in the DSY, so part (b) holds when $d\geq 2$.
Likewise, when $\gamma\leq(d+3)/6$, we have $I=I_{+}+I_{-}\leq 0$, and therefore $\mathbb{E}(\ln R_{\min})<0$, which by Section 3.2 implies non-hyperexplosion, establishing part (c).
Remark 3.16.
The above proof shows that $\mathbb{E}[\ln R_{1}]<0$ when $\gamma<(d+3)/6$, $d\geq 2$. As in the proof of part (c) of Section 3.2, this implies that any (deterministic) ray $s\in\partial\mathbb{T}$ the Markov chain $\{Y_{s|n}\}_{n}$ is non-exploding, in the sense that a.s.
$\tau_{s}=\sum_{j=0}^{\infty}\frac{T_{\bar{s}|j}}{|W_{\bar{s}|j}|^{2\gamma}}=\sum_{j=0}^{\infty}\frac{T_{\bar{s}|j}}{\prod_{k=0}^{j}R_{\max,k}^{2\gamma}}=\infty\,$,
Moreover, by the symmetry property in Section 3.1, $\tau_{s}=\infty$ in the limit case $\gamma=(d+3)/6$ ($d\geq 2$), which includes the standard Navier-Stokes case $d=3$, $\gamma=1$.
Proposition 3.17.
For $d=1$ and $\gamma\in(1/2,3/4)$, the pdf of $R_{\max}=\max\{R_{1},R_{2}\}$ and $R_{\min}=\min\{R_{1},R_{2}\}$ are given by
[…]
Proof.
Assume, without loss of generality, $|\xi|=1$. There are two cases $\xi=1$ or $\xi=-1$. In view of the even parity of $h$, it suffices to consider the case $\xi=1$. Then $R_{1}=|W_{1}|$ and $R_{2}=|1-W_{1}|$. Thus, $R_{\max}\geq 1/2$. For any $r\geq 1/2$,
$\mathbb{P}({{R}_{\max}}\leq r)$ $=$ $\mathbb{P}({{R}_{1}}\leq{{R}_{2}}\leq r)+\mathbb{P}({{R}_{2}}\leq{{R}_{1}}\leq r)$
$=$ $\mathbb{P}({{W}_{1}}\leq 1/2,1-{{W}_{1}}\leq r)+\mathbb{P}({{W}_{1}}\geq 1/2,{{W}_{1}}\leq r)$
$=$ $2\mathbb{P}(1/2\leq{{W}_{1}}\leq r)$
$=$ $2\int\limits_{1/2}^{r}{H(s|1)ds}.$
Therefore, the probability density function of $R_{\max}$ is given by
[…]
which is equivalent to (3.31). Similarly,
$\mathbb{P}(R_{\min}\geq r)=2\mathbb{P}(W_{1}\geq 1/2,|1-W_{1}|\geq r)=2(\mathbb{P}(1/2\leq W_{1}\leq 1-r)+\mathbb{P}(W_{1}\geq 1+r)).$
Taking the derivative of both sides, one gets
[…]
which is equivalent to (3.34). ∎
3.3 Numerical results
Figure 6 is a $d\gamma$-diagram that shows the regions of nonexplosion, explosion, and hyperexplosion described by the criteria in Section 3.2. Although the dimension $d$ has to be an integer, the criteria involve estimates that remain valid for any real values $d>0$. We observe the following:
  1. 1)
    For each fixed $d$, the non-explosion region of $(d,\gamma)$ corresponds to low values of $\gamma$, while hyper-explosion corresponds to high values of $\gamma$. As $\gamma$ increases, there are transitions from non-explosion to explosion (but non-hyperexplosion), and to hyperexplosion.
  2. 2)
    For $d\in\{3,4,5\}$ and $\gamma=1$, the DSY cascade is explosive and non-hyperexplosive. This extends the result of [16, Prop. 12.8] where explosion is proven for $d=3$ and $\gamma=1$. For $d\geq 12$ and $\gamma=1$, the DSY cascade is non-explosive, consistent with the theory; [15, Prop. 12.3].
  3. 3)
    For $d\in\{1,2\}$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$, the DSY cascade is explosive.
  4. 4)
    For $d=1$, the non-hyperexplosion criterion
    $\mathbb{E}[\ln R_{\min}]=\int_{0}^{\infty}g_{R_{\min}}(r)\ln rdr<0$
    infers that the cascade is non-hyperexplosive for $\gamma\in(0.5,\,0.6849]$. Here, $g_{R_{\min}}$ is given by (3.34). With the aid of Mathematica, one can obtain an explicit form of the moment $f(b)=\mathbb{E}[R^{b}]=\int_{0}^{\infty}g_{1,\gamma}(r)r^{b}dr$, where $g_{1,\gamma}$ is the pdf of $R$ given by (3.5), as follows.
    $f(b)=\frac{\Gamma(1-\gamma)^{2}\Gamma(2\gamma-1)^{2}(\Gamma(b+2\gamma-1)\Gamma(-b-2\gamma+2)+\Gamma(b+4\gamma-2)\Gamma(-b-4\gamma+3))}{\sqrt{\pi}\Gamma\left(\frac{3}{2}-2\gamma\right)\Gamma\left(\gamma-\frac{1}{2}\right)^{2}\Gamma(-b-2\gamma+2)\Gamma(b+4\gamma-2)}$
    The hyperexplosion criterion
    $\mathbb{E}[R^{-2\gamma\alpha}]=\int_{0}^{\infty}g_{R_{\min}}(r)r^{-2\gamma\alpha}dr<0.5$
    for some $\alpha\in(0,1-\frac{1}{2\gamma})$ infers that the cascade is hyperexplosive for $\gamma\in[0.7119,\,0.75)$ (Figure 5).
    Figure 5: Graph of $f(b)=\mathbb{E}[R^{b}]$ for $d=1,\,\gamma=0.72$. Its minimum value is approximately $0.4318$, attained at $b\approx-0.16$. The hyperexplosion criterion Section 3.2 (d) is satisfied with the choice $\alpha=-\frac{b}{2\gamma}\approx\frac{1}{9}$.
Figure 6: $d\gamma$-diagram. The horizontal $d$-scale is discretized with stepsize 0.1, and the vertical $\gamma$-scale is discretized with stepsize 0.01. Two different ranges of $d$ are shown: $d\in[1.5,55]$ in the upper figure, and $d\in[1.7,12.2]$ in the lower figure.
4 Majorizing Principle, Integrability of the Solution Process and the Fractional Montgomery-Smith equation
The quadratic form $a\odot_{\xi}b$ defined by (2.8) satisfies the following estimate:
[…]
The proof of (4.1) is elementary and can be found in [18, Lem. 2.1].
To establish the integrability of the solution process $\mathbf{X}$ defined (2.14) it is natural to use the majorization (4.1) of the quadratic form $a\odot_{\xi}b$. In so doing, we establish a natural relation with equation (MS), which will be referred in this paper as the (fractional) Montgomery-Smith equation. The author of [32] introduced (MS) for the case $d=3,\gamma=1$, as a tool to explore the limitation of methods of functional analysis in establishing the well possednes of the Navier-Stokes equations. In the context of $(d,\gamma)$-DSY cascades, the Montgomery-Smith equation plays two roles. The first role is as the governing equation for the probability of no explosion
$\rho(\xi,t)=P_{\xi}(S>t),\quad\rho(\xi,0)=1,$
of the $(d,\gamma)$-DSY cascade, see Section 4. The other role is that it is a tool to obtain a lower bound for $t_{c}$, defined in (2.16) as the time at which the solution process fails to be integrable, see Section 4.
For the first, recall that in Section 3 of the $(d,\gamma)$-DSY cascade, the conditional distribution of the wave vector $W_{v1}$ given $W_{v}=\xi$ is $H(\eta|\xi)=h(\eta)h(\xi-\eta)/(|\xi|^{2\gamma-1}h(\xi))$ where $h$, given by (2.10), is a positive function satisfying the fundamental relation (2.11) $h*h(\xi)=|\xi|^{2\gamma-1}h(\xi)$. We then have,
Proposition 4.1.
Given $(d,\gamma)$, $\rho(\xi,t)$ satisfies
$\rho(\xi,t)=e^{-|\xi|^{2\gamma}t}+\int_{0}^{t}|\xi|^{2\gamma}e^{-|\xi|^{2\gamma}s}\int_{\mathbb{R}^{d}}\rho(\eta,t-s)\rho(\xi-\eta,t-s)H(\eta|\xi)d\eta ds.$ (4.2)
Consequently, $\rho$ is continuous in time $t$.
Proof.
This follows similar lines as the proof of Section 2. Note that $\mathscr{X}(\xi,t)={\mathds{1}_{S_{\xi}>t}}$ satisfies
$\mathscr{X}(\xi,t)=\left\{\begin{array}[]{ll}1,&Y_{\emptyset}\geq t,\\ \mathscr{X}^{(1)}(\eta,t-Y_{\emptyset})\mathscr{X}^{(2)}(\xi-\eta,t-Y_{\emptyset}),\quad&Y_{\emptyset}<t,\end{array}\right$. (4.3)
where $\mathscr{X}^{(1)}$ and $\mathscr{X}^{(2)}$ are two conditionally independent copies on $\mathscr{X}$. Indeed, this follows from the inclusion $Y_{\emptyset}\geq t]\subset[S>t]$ and from the inheritance property of the non-explosion event: ${{\mathds{1}_{S>t}}}{\mathds{1}_{Y_{\emptyset}<t}}={\mathds{1}_{Y_{\emptyset}<t}}\left({\mathds{1}_{{{S}_{{{W}_{1}}}}>t-Y_{\emptyset}}}+{\mathds{1}_{{S_{{{W}_{2}}}}>t-Y_{\emptyset}}}\right)$. Since $|\bar{\mathbf{X}}|\leq 1$, we can take expectation in (4.3) to obtain that $\rho=\mathbb{E}\bar{\mathbf{X}}$ satisfies (4.2). The continuity of $\rho$ follows immediately from (4.2) since $\int_{\mathbb{R}^{3}}\rho(\eta,t-s)\rho(\xi-\eta,t-s)H(\eta|\xi)d\eta\leq 1.$
Note that the constant function $1$ is a trivial solution of (4.2). Consequently, determining whether stochastic explosion occurs, i.e. $\rho(\xi,t)<1$ for some $(\xi,t)$) is equivalent to determining whether (4.2) has other solutions besides the trivial solution for the same initial data $\rho(\xi,0)-1$. This is indeed the case, since as shown in Section 3 there exists values of $(d,\gamma)$ such that the resulting DSY cascade is explosive.
Generalizing the equation introduced in [32], consider the fractional Montgomery-Smith equation in spatial dimension $d\geq 1$ given by
${{\partial}_{t}}v+(-\Delta)^{\gamma}v=(-\Delta)^{1/2}({{v}^{2}})$ (MS)
with initial data $v_{0}(\xi)$. Here, $v(\xi,t)$ and $v_{0}(\xi)=v(\xi,0)$ are scalar valued functions. Note that (MS) has the same scaling property as the fractional Navier-Stokes equations (NSE)d,γ. The Fourier transform of (MS) is given, with $c_{0}=(2\pi)^{-d/2}$, by
$\frac{\partial\hat{v}}{\partial t}+|\xi|^{2\gamma}\hat{v}=c_{0}|\xi|\int_{\mathbb{R}^{d}}\hat{v}(\eta,t)\hat{v}(\xi-\eta,t)d\eta,~~~~\hat{v}(\xi,0)=\hat{v}_{0}(\xi).$
The mild formulation of this equation is
$\hat{v}=e^{-t|\xi|^{2\gamma}}\hat{v}_{0}+c_{0}\int_{0}^{t}|\xi|{{{e}^{-s|\xi{{|}^{2\gamma}}}}\int_{{{\mathbb{R}}^{d}}}{\hat{v}(\eta,t-s)\hat{v}(\xi-\eta,t-s)d\eta ds}}.$ (FMS)
We now relate the explosion problem of the $(d,\gamma)$-DSY cascade with non-uniqueness of solutions of (MS).
Corollary 4.2.
Assume $(d,\gamma)$ are such that the corresponding DSY cascade is explosive and let $\rho(\xi,t)$ denote the probability of no-explosion. Then
$\hat{v}_{1}(\xi)=c_{0}^{-1}h(\xi),~~~\textrm{and}~~~\hat{v}_{\rho}={c_{0}^{-1}h(\xi)}\rho(\xi,t)$
are two solutions of (FMS) with the same initial data $c^{-1}_{0}h(\xi).$
Proof.
We illustrate the proof considering the case $v_{\rho}$ as the case for $v_{1}$ follows similar steps. Rewriting (4.2) in terms of $v_{\rho}$, we get
$\hat{v}_{\rho}(\xi,t)\frac{c_{0}}{h(\xi)}=e^{-t|\xi|^{2\gamma}}+c_{0}^{2}\int_{0}^{t}|\xi|^{2\gamma}e^{-|\xi|^{2\gamma}s}\int_{\mathbb{R}^{d}}\frac{\hat{v}_{\rho}(\eta,t-s)}{h(\eta)}\frac{\hat{v}_{\rho}(\xi-\eta,t-s)}{h(\xi-\eta)}H(\eta|\xi)d\eta ds.$
Recalling the definition of $H(\eta|\xi)$ given in (2.12), we obtain
$\hat{v}_{\rho}(\xi,t)\frac{c_{0}}{h(\xi)}=e^{-t|\xi|^{2\gamma}}+\frac{c_{0}^{2}}{h(\xi)}\int_{0}^{t}|\xi|e^{-|\xi|^{2\gamma}s}\int_{\mathbb{R}^{d}}{\hat{v}_{\rho}(\eta,t-s)}{\hat{v}_{\rho}(\xi-\eta,t-s)}d\eta ds$
which is equivalent to (FMS) with initial data $h(\xi)/c_{0}$.
As noted before the constant $1$ is a solution of (4.2) since the kernel $h(\xi)$ defined in (2.10) satisfies (2.11). ∎
In what follows it will be necessary to introduce the normalized Fourier Montgomery-Smith equation given by
$\psi(\xi,t)={{e}^{-t|\xi{{|}^{2\gamma}}}}{{\psi}_{0}}(\xi)+\int_{0}^{t}{{{e}^{-s|\xi{{|}^{2\gamma}}}}|\xi{{|}^{2\gamma}}\int_{{{\mathbb{R}}^{d}}}{\psi(\eta,t-s)\psi(\xi-\eta,t-s)H(\eta|\xi)d\eta ds}}.$ (nFMS)
[…]
Proposition 4.3.
Let $\chi_{0}:\mathbb{R}^{d}\to\mathbb{C}^{d}$ be a measurable function and let $\psi_{0}=|\chi_{0}|/2$. We have the following statements.
  1. (a)
[…]
  2. (b)
    For each $\xi\neq 0$, $e^{t|\xi|^{2}}\psi(\xi,t)$ is nondecreasing in $t$.
  3. (c)
    (Majorizing principle) Suppose that for some $T\in(0,\infty]$, $\psi<\infty$ a.e. on $\mathbb{R}^{d}\times(0,T)$. Then there exists a subset $D\subset\mathbb{R}^{d}$ with full measure such that both $\psi(\xi,t)=\mathbb{E}_{\xi}\bar{\mathbf{X}}$ and $\mathbb{E}_{\xi}[|{\mathbf{X}}|\mathbbm{1}_{S>t}]$ are finite on $D_{T}=D\times[0,T)$. Moreover, $|\chi|\leq 2\psi$ on $D_{T}$, where $\chi(\xi,t)=\mathbb{E}_{\xi}[{\bf{X}}\mathbbm{1}_{S>t}]$.
Proof.
[…]
$e^{t|\xi|^{2\gamma}}\psi-\psi_{0}(\xi)=\int_{0}^{t}|\xi|^{2\gamma}e^{s|\xi|^{2\gamma}}f(\xi,s)ds$ (4.5)
where $f(\xi,s)=\int\psi(\eta,s)\psi(\xi-\eta,s)H(\eta|\xi)d\eta\geq 0$. The monotonicity of $e^{t|\xi|^{2\gamma}}\psi$ immediately follows.
[…]
$\mathbb{E}_{\xi}[|{\bf{X}}|{\mathds{1}_{S>t}}]\leq 2\mathbb{E}_{\xi}\bar{\mathbf{X}}<\infty$
for any $(\xi,t)\in D_{T}$ and $s\in(t,T)$. ∎
Remark 4.4.
According to Section 4(c), if (FMS) has a mild solution on $(0,\tau)$ then $\mathbb{E}_{\xi}[|{\bf{X}}|\mathbbm{1}_{S>t}]<\infty$ a.e. on $\mathbb{R}^{d}\times(0,\tau)$ which implies the lower bound on the critical time $t_{c}\geq\tau$. The following result (due to Ferreira and Villamizar-Roa [23]) asserts the local-in-time existence and uniqueness of mild solutions to (FMS) for initial data $v_{0}$ in the pseudomeasure-type space $PM^{a}$ defined by (2.3).
Proposition 4.5.
Let $d\geq 2$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$. Let $v_{0}\in PM^{a}$ for some $a\in(d+1-2\gamma,d)$. Then there exists a number $\tau\in(0,\infty)$ such that (FMS) has a unique solution $v\in BC((0,\tau),PM^{a})$. Moreover, if $a=d+1-2\gamma$ and $\|v_{0}\|_{PM^{a}}$ is sufficiently small then one can take $\tau=\infty$.
Proof.
Note that the proof of Theorem 1.2, 1.3, and Remark 1.5(b) of [23] for the special case $a=b$ and $\theta=1$ does not use any special structure of the $\odot$-product other than the fact that $|a\odot_{\xi}b|\leq|a||b|$. Thus, the results hold for the fractional Montgomery-Smith equation. ∎
In [18], where the case $d\geq 3$ and $\gamma=1$ is considered, the authors showed that the solution blows up in finite time for sufficiently large compactly supported initial data. The same result holds for the general case $d\geq 1$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$. More specifically, we have:
Proposition 4.6.
[…]
Proof.
[…]
$\psi(\xi,t)\geq\sum\limits_{n=1}^{\infty}{{{a}^{n}}{{q}_{n}}(\xi,t)}.$ (4.6)
We have $q_{1}(\xi,t)=e^{-|\xi|^{2\gamma}t}\mathbbm{1}_{D}(\xi)$. By conditioning on the first time of branching, one gets
${{q}_{n}}(\xi,t)=\int_{0}^{t}{|\xi|^{2\gamma}{{e}^{-s|\xi|^{2\gamma}}}\int_{{{\mathbb{R}}^{d}}}{\sum\limits_{k=1}^{n-1}{{{q}_{k}}(\eta,t-s){{q}_{n-k}}(\xi-\eta,t-s)H(\eta|\xi)d\eta ds}}}.$ (4.7)
From now on, the proof follows almost verbatim the proof of Proposition 4.14 in [18], which we will leave to the reader. ∎
Remark 4.7.
The majorizing kernel $h:\mathbb{R}\to(0,\infty]$ does not exist in the case $d=\gamma=1$ (see [34]), so our probabilistic model does not apply to this case. Nevertheless, Sverak and Polacik consider the complex Burgers equation ($d=\gamma=1$) and obtain a similar result [35]. Namely, the equation has a global solution $u\in L^{1}(\mathbb{R}\times(0,1))$ provided that $\|\textup{Im}(u_{0})\|_{L^{1}}\leq 2\pi$. Moreover, they show that the upper bound $2\pi$ is sharp in the sense that it is the supremum of the set of $\rho>0$ such that if $\|u_{0}\|_{L_{1}}<\rho$ then (2.20) is globally well-posed.
5 Equations in $\mathbb{R}^{2}$ - Role of Symmetries in Extending Solutions Beyond $t_{c}$.
As pointed out in Section 2, the bilinear operator associated with the fractional Navier-Stokes equations in 2-dimensions simplifies, leading to an essentially scalar operation. In this section we exploit this structure to express the solution process corresponding to these equations as a product of scalars on the corresponding binary tree. This structure is further used to show that $t_{c}$ defined in (2.16) is finite for some specially constructed initial data. Nevertheless, it is possible under some further conditions on the initial data, to extend these solutions beyond this critical time, leading to a global solution of the equations. This extension is possible due to the symmetries of the solution process with respect to tree automorphisms that preserve the DSY structure.
We begin by obtaining an expression of the solution operator that involves a product of scalar quantities determined in terms of the initial data, and the angles between the labels of the $(d,\gamma)-DSY$.
Proposition 5.1.
Let $d=2$ and $\gamma\in(1/2,1)$, and let $\chi_{0}:\mathbb{R}^{2}\to\mathbb{C}^{2}$ be vector field satisfying (2.13). One has
$\mathbf{X}(\xi,t)\mathbbm{1}_{[S>t]}=\left(\frac{i}{2}\right)^{|\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)|}\left(\prod_{v\in\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)}\textup{sign}(W_{v}\times W_{v1})\sin(\theta_{W_{v},W_{v1}}-\theta_{W_{v},W_{v2}})\right)\cdot$
$\left(\prod_{v\in\partial V(\xi,t)}\chi_{0}(W_{v})\cdot e_{W_{v}^{\perp}}\right)e_{\xi^{\perp}}$ (5.1)
Proof.
Let $N(\xi,t)=|\hskip-3.61371pt\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)|$. We will prove by induction on $N(\xi,t)$. If $N(\xi,t)=0$ then $\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)=\emptyset$, $\partial V(\xi,t)=\{\emptyset\}$ and
$\mathbf{X}(\xi,t)=\chi_{0}(\xi,t)=(\chi_{0}(\xi,t)\cdot e_{\xi^{\perp}})e_{\xi^{\perp}}.$
Thus, (5.1) holds for the base case $N(\xi,t)=0$. Suppose that (5.1) holds for any $N(\xi,t)\leq n$, for some $n\geq 0$. For $N(\xi,t)=n+1\geq 1$, we have $Y_{\emptyset}<t$. Thus,
$\mathbf{X}(\xi,t)=\mathbf{X}^{(1)}(W_{1},t-Y_{\emptyset})\odot_{\xi}\mathbf{X}^{(2)}(W_{2},t-Y_{\emptyset}).$ (5.2)
Note that
$\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)$ $=$ $\{\emptyset\}\cup\stackrel{{\scriptstyle\circ}}{{V}}^{(1)}(W_{1},t-Y_{\emptyset})\cup\stackrel{{\scriptstyle\circ}}{{V}}^{(2)}(W_{2},t-Y_{\emptyset})$ (5.3)
$\partial V(\xi,t)$ $=$ $\partial V^{(1)}(W_{1},t-Y_{\emptyset})\cup\partial V^{(2)}(W_{2},t-Y_{\emptyset})$ (5.4)
where $\stackrel{{\scriptstyle\circ}}{{V}}^{(k)}$ and $\partial V^{(k)}$ are the set of internal nodes and the set of leaves of the subtree rooted at $W_{k}$. Since $N^{(1)}(W_{1},t-Y_{\emptyset})+N^{(2)}(W_{2},t-Y_{\emptyset})=N(\xi,t)-1=n$, one deduces by the induction hypothesis that for $k\in\{1,2\}$,
$\mathbf{X}^{(k)}(W_{k},t-Y_{\emptyset})$ $=$ $\left(\frac{i}{2}\right)^{N^{(k)}(W_{k},t-Y_{\emptyset})}\cdot$
$\left(\prod_{v\in\stackrel{{\scriptstyle\circ}}{{V}}^{(k)}(W_{k},t-Y_{\emptyset})}\textup{sign}(W_{v}\times W_{v1})\sin(\theta_{W_{v},W_{v1}}-\theta_{W_{v},W_{v2}})\right)\cdot$
$\left(\prod_{v\in\partial V^{(k)}(W_{k},t-Y_{\emptyset})}\chi_{0}(W_{v})\cdot e_{W_{v}^{\perp}}\right)e_{W_{k}^{\perp}}$
Applying (5.2)-(5.4) and the following identity (proven in Section 2)
$e_{W_{1}^{\perp}}\odot_{\xi}e_{W_{2}}^{\perp}=\frac{i}{2}\text{sign}(\xi\times W_{1})\sin(\theta_{\xi,W_{1}}-\theta_{\xi,W_{2}})e_{\xi^{\perp}},$
one obtains (5.1). ∎
Proposition 5.2.
For any $v\in\mathbb{T}$, let $\Sigma_{v}=\textup{sign}(W_{v}\times W_{v1})$ and
$\Theta_{v}=(\theta_{W_{v},W_{v1}},\theta_{W_{v},W_{v2}})\in\triangle:=\{(\phi_{1},\phi_{2}):\ \phi_{1},\phi_{2}\geq 0,\ \phi_{1}+\phi_{2}\leq\pi\}.$
Then
  1. (a)
    $\{\Sigma_{v}\}_{v\in\mathbb{T}}$ is an i.i.d. family of Bernoulli random variables with parameter $1/2$.
  2. (b)
    $\{\Theta_{v}\}_{v\in\mathbb{T}}$ is an i.i.d. family of random vectors whose joint pdf is given by (3.9).
  3. (c)
    $\{\Sigma_{v}\}_{v\in\mathbb{T}}$ and $\{\Theta_{v}\}_{v\in\mathbb{T}}$ are independent each other.
Proof.
To prove (a), notice that the vector $W_{v}$ partitions the plane into two halves
$A$ $=$ $\{\eta\in\mathbb{R}^{2}:\ W_{v}\times\eta=W_{v}^{\perp}\cdot\eta>0\},$
$B$ $=$ $\{\eta\in\mathbb{R}^{2}:\ W_{v}\times\eta=W_{v}^{\perp}\cdot\eta<0\}.$
We have
$\mathbb{P}_{\xi}(\Sigma_{v}=1)=\mathbb{P}_{\xi}(W_{v1}\in A)$ $=$ $\mathbb{E}_{\xi}[\mathbb{P}_{\xi}(W_{v1}\in A|W_{v})]=\mathbb{E}_{\xi}\left[\int_{A}H(\eta|W_{v})d\eta\right].$
Using the change of variable $\zeta=W_{v}-\eta$ and noting that $H(\eta,W_{v})=H(\zeta,W_{v})$, we have
$\mathbb{P}_{\xi}(\Sigma_{v}=1)=\mathbb{E}_{\xi}\left[\int_{B}H(\zeta|W_{v})d\eta\right]=\mathbb{E}_{\xi}\left[\int_{B}H(\eta|W_{v})d\eta\right]=\mathbb{P}_{\xi}(W_{v1}\in B)=\mathbb{P}_{\xi}(\Sigma_{v}=-1).$
Therefore, $\Sigma_{v}\sim\text{Bernoulli}(1/2)$.
Part (b) and (c) are obvious each vector $\Theta_{v}$ has a joint pdf given by (3.9). ∎
Thanks to Section 5, we can write
$\mathbf{Z}(\xi,t):=\mathbf{X}(\xi,t)\mathbbm{1}_{S>t}=\phi_{\xi,t}(\{\Sigma_{v}\}_{v\in\mathbb{T}},\{\Theta_{v}\}_{v\in\mathbb{T}}).$ (5.5)
Consider the transformation $(\{\Sigma_{v}\}_{v\in\mathbb{T}},\{\Theta_{v}\}_{v\in\mathbb{T}})\mapsto(\{\tilde{\Sigma}_{v}\}_{v\in\mathbb{T}},\{\tilde{\Theta}_{v}\}_{v\in\mathbb{T}})$ in which
$\tilde{\Theta}_{v}$ $=$ $\Theta_{v}\ \ \ \forall v\in\mathbb{T},$ (5.6)
[…]
This transformation induces a transformation $\{W_{v}\}_{v\in\mathbb{T}}\to\{\tilde{W}_{v}\}_{v\in\mathbb{T}}$ in which $\tilde{W}_{\emptyset}=W_{\emptyset}=\xi$ and $|\tilde{W}_{v}|=|W_{v}|$ for all $v\in\mathbb{T}$. The DSY cascade $\{Y_{v}=|W_{v}|^{-2\gamma}T_{v}\}_{v\in\mathbb{T}}$ is then transformed into $\{\tilde{Y}_{v}=|\tilde{W}_{v}|^{-2\gamma}T_{v}=Y_{v}\}_{v\in\mathbb{T}}$. The explosion time is transformed into
$\tilde{S}=\underset{n\geq 0}{\mathop{\sup}}\,\underset{|v|=n}{\mathop{\min}}\,\sum\limits_{j=0}^{n}\tilde{Y}_{v|j}=\underset{n\geq 0}{\mathop{\sup}}\,\underset{|v|=n}{\mathop{\min}}\,\sum\limits_{j=0}^{n}{Y}_{v|j}=S.$
To use Section 5 to explore the possibilities for stochastic cancellation we consider the well-known simple vortex model, see e.g. [30, p 48]. The stream function corresponding to the initial velocity is radially symmetric, and thus, in this case $\hat{u}_{0}(x)=h(|\xi|)(-\xi_{2},\xi_{1})^{t}$. We will refer to this data as a vortex and note that after scaling, $\chi_{0}(\xi)=\rho(|\xi|)e_{\xi^{\perp}}$ for an appropriate complex valued function $\rho$. Section 5 reveals a stochastic asymmetry that results in substantial cancellations in the proof of Section 5 for this type of initial data.
Corollary 5.3.
Let $\tilde{\mathbf{X}}(\xi,t)=\phi_{\xi,t}(\{\tilde{\Sigma}_{v}\}_{v\in\mathbb{T}},\{\tilde{\Theta}_{v}\}_{v\in\mathbb{T}})$ where $\{\tilde{\Sigma}_{v}\}_{v\in\mathbb{T}}$ and $\{\tilde{\Theta}_{v}\}_{v\in\mathbb{T}}$ are given by (5.6)-(5.9). Suppose the vector field $\chi_{0}$ is a vortex. Then
  1. (a)
    $\mathbf{X}\mathbbm{1}_{T_{\emptyset}<t<S}=-\tilde{\mathbf{X}}\mathbbm{1}_{T_{\emptyset}<t<{S}}$.
  2. (b)
    If $\mathbb{E}[|\mathbf{X}(\xi,t)|\mathbbm{1}_{T_{\emptyset}<t<S}]<\infty$ then $\mathbb{E}[\mathbf{X}(\xi,t)\mathbbm{1}_{T_{\emptyset}<t<S}]=0$.
Proof.
To show Part (a), notice from (5.1) that
$\mathbf{X}\mathbbm{1}_{S>t}=\left(\frac{i}{2}\right)^{|\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)|}\left(\prod_{v\in\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)}\Sigma_{v}\sin(\Theta_{v1}-\Theta_{v2})\right)\left(\prod_{v\in\partial V(\xi,t)}\rho(|W_{v}|)\right)e_{\xi^{\perp}}$ (5.10)
where $\rho(\eta)=\chi_{0}(\eta)\cdot e_{\eta^{\perp}}$. On the event $T_{\emptyset}<t<S]$, one has $\emptyset\in\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)$. Using the fact that $\tilde{\Sigma}_{\emptyset}=-\Sigma_{\emptyset}$, one has
$\tilde{\mathbf{X}}\mathbbm{1}_{[T_{\emptyset}<t<S]}=\left(\frac{i}{2}\right)^{|\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)|}\left(\prod_{v\in\stackrel{{\scriptstyle\circ}}{{V}}(\xi,t)}\tilde{\Sigma}_{v}\sin(\Theta_{v1}-\Theta_{v2})\right)\left(\prod_{v\in\partial V(\xi,t)}\rho(|W_{v}|)\right)e_{\xi^{\perp}}=-\mathbf{X}\mathbbm{1}_{[T_{\emptyset}<t<S]}.$ (5.11)
To show (b), notice that the family $\{\tilde{\Sigma}_{v}\}_{v\in\mathbb{T}}$ has the same distribution as $\{{\Sigma}_{v}\}_{v\in\mathbb{T}}$. Assuming the integrability condition $\mathbb{E}[|\mathbf{X}(\xi,t)|\mathbbm{1}_{T_{\emptyset}<t<S}]<\infty$, one has
$\mathbb{E}[\mathbf{X}(\xi,t)\mathbbm{1}_{T_{\emptyset}<t<S}]=\mathbb{E}[\tilde{\mathbf{X}}(\xi,t)\mathbbm{1}_{T_{\emptyset}<t<S}].$
On the other hand, by Part (a) one has $\mathbb{E}[\mathbf{X}(\xi,t)\mathbbm{1}_{T_{\emptyset}<t<S}]=-\mathbb{E}[\tilde{\mathbf{X}}(\xi,t)\mathbbm{1}_{T_{\emptyset}<t<S}]$. Therefore, $\mathbb{E}[\mathbf{X}(\xi,t)\mathbbm{1}_{T_{\emptyset}<t<S}]=0$. ∎
The radial symmetry of the initial data and the invariance of the DSY process leads to the following consequence, also noted in [30, p 48] in the physical space setting.
Corollary 5.4.
Suppose $u_{0}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ is a vortex. Then $\textup{(NSE)}_{d,\gamma}$ has a global solution $u$ whose Fourier transform is given by $\hat{u}(\xi,t)=\hat{u}_{0}(\xi)e^{-|\xi|^{2\gamma}t}$.
Proof.
Define
[…]
Then clearly $\hat{u}(\xi,t)=\mathbb{E}\mathbf{X}^{*}(\xi,t)$ is a solution for $t<t_{c}(\hat{u}_{0}(\xi))$. In view of Corollary 5, one has
${\mathbf{X}}^{*}(\xi,t)=\hat{u}_{0}(\xi){\mathds{1}_{[T_{\emptyset}>t]}}.$
Thus, $u(\xi,t)=\hat{u}_{0}(\xi)e^{-|\xi|^{2\gamma}t}$ is a solution for $t<t_{c}(\hat{u}_{0}(\xi))$. However, one may readily check that the validity of this mild solution extends to all times $t$ beyond $t_{c}(\hat{u}_{0}(\xi))$ as well. ∎
Remark 5.5.
Section 5 can also be proven analytically in Fourier space. For convenience, the proof is provided in Section 6. However, our approach highlights the role symmetries of DSY cascades play in providing cancellations necessary to prove global existence of solutions of $\textup{(NSE)}_{d,\gamma}$.
Proposition 5.6.
Let $t_{c}$ be defined by (2.16). Denote $f(\xi,t)=\mathbb{E}_{\xi}[|\mathbf{X}(\xi,t)|\mathbbm{1}_{S>t}]\in[0,\infty]$. Then:
  1. (a)
    The function $v(\xi,t)=\frac{h(\xi)}{c_{0}}f(\xi,t)$ satisfies
    $v(\xi,t)=v_{0}(\xi)e^{-|\xi|^{2\gamma}t}+\frac{c_{0}}{2}\int_{0}^{t}\int_{\mathbb{R}^{2}}|\xi|e^{-|\xi|^{2\gamma}s}v(\eta,t-s)v(\xi-\eta,t-s)|\sin(\theta_{\xi,\eta}-\theta_{\xi,\xi-\eta})|d\eta ds$ (5.12)
[…]
  2. (b)
    If $t_{c}>0$ then $f(\xi,t)<\infty$ for almost every $(\xi,t)\in\mathbb{R}^{2}\times(0,t_{c})$.
  3. (c)
    If
    $\exists A\subset\mathbb{R}^{2}:m(A)>0\text{~~and~~}\chi_{0}(\xi)\neq 0\text{~~a.e.~~}\xi\in A\cup(-A)$ (5.13)
[…]
  4. (d)
    If $t_{c}<\infty$ and (5.13) is satisfied then $f(\xi,t)=\infty$ for all $(\xi,t)\in\mathbb{R}^{2}\times(t_{c},\infty)$.
Proof.
(a) According to (2.14) and Section 5,
[…]
Let $v(\xi,t)=\frac{h(\xi)}{c_{0}}f(\xi,t)$. Taking the expectation of (5.14), one obtains (5.12).
(b) Let
$E$ $=$ $\{(\xi,t)\in\mathbb{R}^{2}\times(0,t_{c}):f(\xi,t)=\infty\},$
$E_{t}$ $=$ $\{\xi\in\mathbb{R}^{2}:f(\xi,t)=\infty\}.$ (5.15)
By the monotonicity of $e^{|\xi|^{2\gamma}t}f(\xi,t)$ as proven in Part (a), $\{E_{t}\}$ is an increasing family. That is, $E_{t}\subset E_{s}$ if $t<s$. We have
$m(E)=\int_{0}^{t_{c}}\int_{\mathbb{R}^{2}}\mathbbm{1}_{E}(\xi,t)d\xi dt=\int_{0}^{t_{c}}m(E_{t})dt.$
Suppose by contradiction that $m(E)>0$. Then there exists $t_{0}\in(0,t_{c})$ such that $m(E_{t_{0}})>0$. The monotonicity of $\{E_{t}\}$ implies $m(E_{s})>0$ for all $t_{0}\leq s<t_{c}$. This contradicts the definition of $t_{c}$.
(c) Assume (5.13). Since $A$ is of positive measure, there exists a number $M>0$ such that $B=\{\xi\in A\cup(-A):|\xi|\leq M\}$ is of positive measure. For any $\zeta\in\mathbb{R}^{2}$, denote
$\zeta+B=\{\zeta+\xi:\xi\in B\}.$
For each $r>0$, let $D_{r}=\{\xi\in\mathbb{R}^{2}:|\xi|<r\}$. By the continuity of the translation operator in $L^{1}(\mathbb{R}^{2})$ (see e.g. [36, Exer. 12, p. 132]),
$\lim_{\zeta\to 0}\int_{\mathbb{R}^{2}}|\mathbbm{1}_{\zeta+B}(\xi)-\mathbbm{1}_{B}(\xi)|d\xi=0.$
Equivalently, $m((\zeta+B)\backslash B)+m(B\backslash(\zeta+B))\to 0$ as $\zeta\to 0$. This implies $m(B\cup(\zeta+B))=m(B)+m((\zeta+B)\backslash B)\to m(B)$ as $\zeta\to 0$. Hence,
$m(B\cap(\zeta+B))=m(B)+m(\zeta+B)-m(B\cup(\zeta+B))=2m(B)-m(B\cup(\zeta+B))\to m(B)$
as $\zeta\to 0$. Thus, there exists $\epsilon>0$ such that $m(B\cap(\zeta+B))>0$ for all $\zeta\in\mathbb{R}^{2},|\zeta|<\epsilon$. Let
$\tilde{E}=\{\xi\in\mathbb{R}^{2}:v(\xi,t)>0\ \ \forall\,t>0\}.$
To show that $\tilde{E}=\mathbb{R}^{2}$, we will show by induction on $n\in\mathbb{N}$ that $D_{n\epsilon}\subset\tilde{E}$.
For any $\xi\in D_{\epsilon}$ and $t>0$, (5.18) implies
$v(\xi,t)\geq\frac{c_{0}}{2}\int_{0}^{t}\int_{B\cap(\xi+B)}|\xi|e^{-|\xi|^{2\gamma}s}v(\eta,t-s)v(\xi-\eta,t-s)|\sin(\theta_{\xi,\eta}-\theta_{\xi,\xi-\eta})|d\eta ds$ (5.16)
For any $\eta\in B\cap(\xi+B)$, we have $\xi-\eta\in B\cap(\xi+B)$ and thus, $v_{0}(\eta)>0$ and $v_{0}(\xi-\eta)>0$. Together with the monotonicity in $t$ of $e^{|\xi|^{2\gamma}t}f(\xi,t)$ as proven in Part (a), we have $v(\eta,s)>0$ and $v(\xi-\eta,s)>0$ for any $s>0$. Then (5.16) implies that $v(\xi,t)>0$, which implies $D_{\epsilon}\subset\tilde{E}$.
Now suppose that $D_{n\epsilon}\subset\tilde{E}$ for some $n\in\mathbb{N}$. Let $\xi\in D_{(n+1)\epsilon}$. Note that the set $D=D_{n\epsilon}\cap(\xi+D_{\epsilon})$ is of positive measure. By (5.18),
$v(\xi,t)\geq\frac{c_{0}}{2}\int_{0}^{t}\int_{D}|\xi|e^{-|\xi|^{2\gamma}s}v(\eta,t-s)v(\xi-\eta,t-s)|\sin(\theta_{\xi,\eta}-\theta_{\xi,\xi-\eta})|d\eta ds.$ (5.17)
For any $\eta\in D$, one has $\eta\in(\xi+D_{\epsilon})\subset E$ and thus, $\xi-\eta\in D_{\epsilon}\subset\tilde{E}$. Hence, $v(\eta,s)>0$ and $v(\xi-\eta,s)>0$ for any $s>0$. By (5.17), $v(\xi,t)>0$. We have shown that $D_{(n+1)\epsilon}\subset\tilde{E}$.
(d) Suppose that $t_{c}<\infty$ and that (5.13) is satisfied. Fix $t>t_{c}$ and let $\tau=(t+t_{c})/2$. The family of sets $\{E_{s}\}$ defined by (5.15) satisfies $E_{\tau}\subset E_{s}$ for any $\tau\leq s\leq t$ and $m(E_{\tau})>0$. On the other hand, by Part (c), one has $v(\zeta,s)>0$ for any $\zeta\in\mathbb{R}^{2}$ and $s>0$. Hence, for any $\xi\in\mathbb{R}^{2}$ and $\eta\in E_{\tau}$, we have $v(\eta,s)=\infty$ and $v(\xi-\eta,s)>0$ for any $\tau\leq s\leq t$. Consequently,
$v(\xi,t)\geq\frac{c_{0}}{2}\int_{0}^{t-\tau}\int_{E_{\tau}}|\xi|e^{-|\xi|^{2\gamma}s}v(\eta,t-s)v(\xi-\eta,t-s)|\sin(\theta_{\xi,\eta}-\theta_{\xi,\xi-\eta})|d\eta ds=\infty.$
Proposition 5.7.
Let $d=2$ and $\gamma\in(1/2,1)$, and let $u_{0}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ be a divergence-free vector field. Suppose
$M=\inf_{4\leq|\xi|\leq 7}|\hat{u}_{0}(\xi)|>500\cdot 7^{2\gamma}.$
Then the set
$D=\{(\xi,t)\in\mathbb{R}^{2}\times(0,T_{*}):\ \mathbb{E}[|\mathbf{X}(\xi,t)|\mathbbm{1}_{[S>t]}]=\infty\}$
where
$T_{*}=-7^{-2\gamma}\ln\left(1-\frac{500\cdot 7^{2\gamma}}{M}\right)$
is of positive measure.
Proof.
Let $\chi_{0}(\xi)=c_{0}\frac{\hat{u}_{0}(\xi)}{h(\xi)}$ and $v(\xi,t)=\frac{h(\xi)}{c_{0}}\mathbb{E}[|\mathbf{X}(\xi,t)|\mathbbm{1}_{[S>t]}]$ where $c_{0}=(2\pi)^{-1}$. Taking the expectation of (5.14), one gets that for all $t\in[0,T^{*}]$,
$v(\xi,t)=v_{0}(\xi)e^{-|\xi|^{2\gamma}t}+\frac{c_{0}}{2}\int_{0}^{t}\int_{\mathbb{R}^{2}}|\xi|e^{-|\xi|^{2\gamma}s}v(\eta,t-s)v(\xi-\eta,t-s)|\sin(\theta_{\xi,\eta}-\theta_{\xi,\xi-\eta})|d\eta ds$ (5.18)
where $v_{0}(\xi)=|\hat{u}_{0}(\xi)|\geq M$ for $4\leq|\xi|\leq 7$. Suppose by contradiction that $D$ is of measure zero. Then $v(\xi,t)<\infty$ for almost every $(\xi,t)\in\mathbb{R}^{2}\times(0,T_{*})$. It is easy to check that $e^{|\xi|^{2\gamma}t}v(\xi,t)$ is an increasing function in $t$. Hence, for each $t\in(0,T_{*})$,
$v(\xi,t)<\infty,\ \ \ \text{a.e.}\ \xi\in\mathbb{R}^{2}.$ (5.19)
Let
$A_{\xi}=\{\eta\in\mathbb{R}^{2}:\ 4\leq|\eta|\leq 5,\ 6\leq|\xi-\eta|\leq 7\}.$
Then for any $\xi$ such that $4\leq|\xi|\leq 7$, one has
$v(\xi,t)\geq e^{-7^{2\gamma}t}M+\int_{0}^{t}\int_{A_{\xi}}\frac{1}{\pi}e^{-7^{2\gamma}s}v(\eta,t-s)v(\xi-\eta,t-s)|\sin(\theta_{\xi,\eta}-\theta_{\xi,\xi-\eta})|d\eta ds$ (5.20)
Let $\beta=7^{2\gamma}$. Multiplying both sides of (5.20) by $e^{\beta t}$, we get
$e^{\beta t}v(\xi,t)\geq M+\int_{0}^{t}\int_{A_{\xi}}\frac{1}{\pi}e^{\beta(t-s)}v(\eta,t-s)v(\xi-\eta,t-s)|\sin(\theta_{\xi,\eta}-\theta_{\xi,\xi-\eta})|d\eta ds$ (5.21)
for any $\xi$ such that $4\leq|\xi|\leq 7$. Next, we find lower bounds for $|A_{\xi}|$ and $|\sin(\theta_{\xi,\eta}-\theta_{\xi,\xi-\eta})|$ provided that
$4\leq|\xi|\leq 7,\ 4\leq|\eta|\leq 5,\ 6\leq|\xi-\eta|\leq 7.$
Estimate of $|\sin(\theta_{\xi,\eta}-\theta_{\xi,\xi-\eta})|$:
Let $\phi_{1}=\sin\theta_{\xi,\eta}\in[0,\pi]$, $\phi_{2}=\sin\theta_{\xi,\xi-\eta}\in[0,\pi]$, $a=|\eta|$, $b=|\xi-\eta|$, $c=|\xi|$. Label points $A,B,C,D$ as in Figure 7.
Figure 7:
By the triangle inequality, $DC>AC-AD=|\xi-\eta|-|\eta|\geq 6-5=1$. By the Sine Law in triangle $ADC$,
$|\sin(\phi_{1}-\phi_{2})|=\frac{2\text{area}(ADC)}{ab}\geq\frac{2\text{area}(ADC)}{5\times 7}=\frac{2\text{area}(ADC)}{35}.$
On the other hand,
$\frac{\text{area}(ADC)}{\text{area}(ABC)}=\frac{DC}{BC}\geq\frac{1}{7}.$
Combining the two above inequalities, one gets
$|\sin(\phi_{1}-\phi_{2})|\geq\frac{2}{245}\text{area}(ABC)$
By Heron’s formula,
$\text{area}(ABC)=\frac{1}{4}\sqrt{(a+b+c)(a+b-c)(a+c-b)(b+c-a)}.$
Under the conditions $4\leq a\leq 5$, $6\leq b\leq 7$, $4\leq c\leq 7$, we have
$\text{area}(ABC)\geq\frac{1}{4}\sqrt{(4+6+4)(4+6-7)(4+4-7)(6+4-5)}=\frac{\sqrt{210}}{4}.$
Thus,
$|\sin(\phi_{1}-\phi_{2})|\geq\frac{2}{245}\frac{\sqrt{210}}{4}=\frac{\sqrt{210}}{490}>\frac{1}{40}.$ (5.22)
Estimate of $|A_{\xi}|$:
For the purpose of estimating $|A_{\xi}|$, we can assume that $\xi$ has coordinates $(c=|\xi|,0)$. Label the points $O(0,0)$, $F(c,0)$, $E\left(-\frac{12}{c}+\frac{c}{2},\sqrt{20-\left(-\frac{12}{c}+\frac{c}{2}\right)^{2}}\right)$ as in Figure 8.
Figure 8:
Then
$OE$ $=$ $\sqrt{\left(-\frac{12}{c}+\frac{c}{2}\right)^{2}+20-\left(-\frac{12}{c}+\frac{c}{2}\right)^{2}}=\sqrt{20},$
$EF$ $=$ $\sqrt{\left(-\frac{12}{c}-\frac{c}{2}\right)^{2}+20-\left(-\frac{12}{c}+\frac{c}{2}\right)^{2}}=\sqrt{44}.$
For any $\eta\in B_{3/10}(E)$,
$|\eta|$ $\leq$ $OE+|\overrightarrow{OE}-\eta|\leq\sqrt{20}+\frac{3}{10}<5,$
$|\eta|$ $\geq$ $OE-|\overrightarrow{OE}-\eta|\leq\sqrt{20}-\frac{3}{10}>4,$
$|\xi-\eta|$ $\leq$ $EF+|\overrightarrow{OE}-\eta|\leq\sqrt{44}+\frac{3}{10}<7,$
$|\xi-\eta|$ $\geq$ $EF-|\overrightarrow{OE}-\eta|\leq\sqrt{44}-\frac{3}{10}>6.$
Therefore, $B_{3/10}(E)\subset A_{\xi}$. Consequently,
$|A_{\xi}|\geq|B_{3/10}|=\pi\left(\frac{3}{10}\right)^{2}=\frac{9\pi}{100}.$ (5.23)
Substituting the estimate (5.22) into (5.21), we get
$e^{\beta t}v(\xi,t)\geq M+\frac{1}{40\pi}\int_{0}^{t}\int_{A_{\xi}}e^{\beta(t-s)}v(\eta,t-s)v(\xi-\eta,t-s)|\sin(\theta_{\xi,\eta}-\theta_{\xi,\xi-\eta})|d\eta ds.$
Let $q(t)=\inf_{4\leq|\xi|\leq 7}e^{\beta t}v(\xi,t)$. Because of (5.19), $q(t)<\infty$ for all $t\in(0,T_{*})$. The inequality above together with (5.23) implies
$q(t)$ $\geq$ $M+\frac{1}{40}\int_{0}^{t}\int_{A_{\xi}}e^{-\beta(t-s)}q^{2}(t-s)d\eta ds$
$=$ $M+\frac{1}{40\pi}|A_{\xi}|\int_{0}^{t}e^{-\beta(t-s)}q^{2}(t-s)ds$
$\geq$ $M+\frac{1}{450}\int_{0}^{t}e^{-\beta s}q^{2}(s)ds$
Then $q(t)\geq p(t)$, where $p(t)$ is the solution to the integral equation
$p(t)=M+\frac{1}{450}\int_{0}^{t}e^{-\beta s}p^{2}(s)ds.$
Differentiating both sides, we get $p^{\prime}=\frac{1}{450}e^{-\beta t}p^{2}$, which has an explicit solution
$p(t)=\frac{1}{\frac{1}{M}-\frac{1}{450\beta}(1-e^{-\beta t})}.$
This function blows up at
$\tau=-\frac{1}{\beta}\ln\left(1-\frac{450\beta}{M}\right).$
Notice that $\tau<T_{*}$. For any $t<\tau$, we have
$q(\tau)$ $\geq$ $M+\frac{1}{450}\int_{0}^{\tau}e^{-\beta s}q^{2}(s)ds$
$\geq$ $M+\frac{1}{450}\int_{0}^{t}e^{-\beta s}q^{2}(s)ds$
$\geq$ $M+\frac{1}{450}\int_{0}^{t}e^{-\beta s}p^{2}(s)ds$
$=$ $p(t).$
Thus, $q(\tau)\geq\limsup_{t\to\tau^{-}}p(t)=\infty$. This is a contradiction. ∎ The fact that the solution process $\mathbf{X}(\xi,t)$ fails to be integrable after finite time for sufficiently large initial condition does not necessarily imply that the mild solution $u$ to $\textup{(NSE)}_{d,\gamma}$ blows up after finite time. It simply implies that the mean-flow representation of the scaled Fourier transform $\chi=c_{0}\frac{\hat{u}}{h}$ ceases to be valid after finite time. For example, if the initial velocity field $u_{0}$ is a vortex, i.e. the stream function is radially symmetric, then the flow evolves but remains a vortex flow (radial eddy) for all time. See e.g. [30, Example 2.2, p. 48], [5, Sec. 4.5], [37, Chap. 13]. Here, a function $f$ on $\mathbb{R}^{2}$ is said to be radially symmetric if $f(x)$ only depends on $|x|$. The same result holds for the fractional Navier-Stokes equations (Section 6).
If $u$ satisfies $\textup{(NSE)}_{d,\gamma}$ then the Fourier transform $v=\hat{u}$ satisfies the equation:
$\textup{(FNSE)}_{d,\gamma}:\ \ \ v=F_{1}(v_{0})+F_{2}(v,v)$ (5.24)
where $v_{0}=\hat{u}_{0}$,
$F_{1}(f)=f(\xi)e^{-|\xi|^{2\gamma}t},\ \ \ F_{2}(f,g)=\int_{0}^{t}|\xi|e^{-|\xi|^{2\gamma}s}\int_{\mathbb{R}^{d}}f(\eta,t-s)\odot_{\xi}g(\xi-\eta,t-s)d\eta ds$
A mild solution to (5.24) is defined as the limit of the Picard iteration
$v^{(n+1)}=F_{1}(v_{0})+F_{2}(v^{(n)},v^{(n)})$ (5.25)
where $v^{(0)}(\xi,t)\equiv 0$.
6 Appendix: Analytical Proof of Section 5
Proposition 6.1.
Suppose $u_{0}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ is a vortex. Then $\textup{(NSE)}_{d,\gamma}$ has a global solution $u$ whose Fourier transform is given by $\hat{u}(\xi,t)=F_{1}(\hat{u}_{0})=\hat{u}_{0}(\xi)e^{-|\xi|^{2\gamma}t}$.
Proof.
Let $\psi$ be a stream function, i.e. $\psi$ satisfies $u_{0}=\nabla^{\perp}\psi=(\psi_{y},-\psi_{x})$. Since $\psi$ is radially symmetric, so is $\hat{\psi}$. One can easily check that $\hat{u}_{0}(\xi)=(i\xi_{2},-i\xi_{1})\hat{\psi}(\xi)$. Thus, $\hat{u}_{0}(\xi)\cdot e_{\xi^{\perp}}=-i|\xi|\hat{\psi}(\xi)$ is also radially symmetric.
Let $v_{0}=\hat{u}_{0}$. We will show that $F_{2}(v^{(n)},v^{(n)})=0$, where $v^{(n)}$ is the sequence defined by (5.25). It suffices to show that $F_{2}(f,f)=0$ if $f=f(\xi,t)$ is a function such that $f(\xi,t)\cdot e_{\xi^{\perp}}$ is radially symmetric in $\xi$. Let $f$ be such a function. By Section 2,
$F_{2}(f,f)(\xi,t)=\frac{i}{2}\int_{0}^{t}|\xi|e^{-|\xi|^{2\gamma}s}\int_{\mathbb{R}^{2}}\text{sign}(\xi\times\eta)g(\eta,t-s)g(\zeta,t-s)\sin(\theta_{\xi,\eta}-\theta_{\xi,\zeta})d\eta ds$
where $\zeta=\xi-\eta$ and $g(\xi,t)=f(\xi,t)\cdot e_{\xi^{\perp}}$. Decompose the integral over $\mathbb{R}^{2}$ into an integral over $\{\eta:\xi\times\eta>0\}$ and an integral over $\{\eta:\xi\times\eta<0\}$ (see Figure 1).
Figure 9:
Then use the mirror reflection with respect to vector $\xi=(\xi_{1},\xi_{2})$, whose matrix representation is
$R_{\xi}=\frac{1}{|\xi|^{2}}\left[\begin{matrix}\xi_{1}^{2}-\xi_{2}^{2}&2{{\xi}_{1}}{{\xi}_{2}}\\ 2{{\xi}_{1}}{{\xi}_{2}}&\xi_{2}^{2}-\xi_{1}^{2}\\ \end{matrix}\right],$
as a change of variable, one obtains
$F(f,f)(\xi,t)=i\int_{0}^{t}|\xi|e^{-|\xi|^{2\gamma}s}\int_{\xi\times\eta>0}g(\eta,t-s)g(\zeta,t-s)\sin(\theta_{\xi,\eta}-\theta_{\xi,\zeta})d\eta ds.$
Let $L\eta=\xi+R_{\xi^{\perp}}\eta$ (see Figure 9). Note that
$|L\eta|=|R_{\xi^{\perp}}(\eta-\xi)|=|\eta-\xi|=|\zeta|,$
$|\xi-L\eta|=|R_{\xi^{\perp}}\eta|=|\eta|,$
$\theta_{\xi,L\eta}=\theta_{\xi,\zeta},\ \ \ \ \theta_{\xi,\xi-L\eta}=\theta_{\xi,\eta}.$
Using the change of variables $\eta\to L\eta$ and the fact that $g$ is radially symmetric, one gets
$\int_{\xi\times\eta>0}g(\eta,t-s)g(\zeta,t-s)\sin(\theta_{\xi,\eta}-\theta_{\xi,\zeta})d\eta=-\int_{\xi\times\eta>0}g(\eta,t-s)g(\zeta,t-s)\sin(\theta_{\xi,\eta}-\theta_{\xi,\zeta})d\eta$
Therefore, $F_{2}(f,f)=0$. ∎